### v1

# Low-Latency ML Inference by Grouping Correlated Data Objects and Computation

Abstract.
ML inference workflows often require low latency and high throughput, yet we lack good options for addressing this need. Techniques that reduce latency in other streaming settings (such as caching and optimization-driven scheduling) are of limited value because ML data dependencies are often very large and can change dramatically depending on the triggering event. In this work, we propose a novel correlation grouping mechanism that makes it easier for developers to express application-specific data access correlations, enabling coordinated management of data objects in server clusters hosting streaming inference tasks. Experiments based on a latency-sensitive ML-based application confirm the limitations of standard techniques while showing that our solution yields dramatically better performance. The proposed mechanism is able to maintain significantly lower and more consistent latency, achieves higher node utilization as workload and scale-out increase, and yet requires only minor changes to the code implementing the application.
conference: Preprint; November 2023; arXiv
1. Introduction
Latency-sensitive AI-based applications are increasingly common (Chen et al., 2021). For instance, in edge intelligence applications, devices pass captured data through Machine Learning (ML) inference pipelines (Deng et al., 2020). Minimizing latency is the highest priority, although high throughput and efficient platform utilization remain important as secondary goals (Crankshaw et al., 2017). Our work starts with the observation that when these pipelines are deployed onto standard Stream Processing (SP) platforms (Isah et al., 2019), a mismatch occurs. SP systems need low latency but are typically optimized for highly parallel stateless computations and data transformations. ML inference is stateful and often context-sensitive, resulting in data dependencies that can defeat the assumptions made in SP schedulers (Shaowang et al., 2021).
The issue is further complicated by unpredictability. ML models depend on pretrained parameter vectors that can be large. For example, a trajectory-computing task for mobile “agents” in a traffic scene might have one set of models trained for cars, another for bicycles, a third for pedestrians, etc. Among the vehicle models there could be one specialized for taxis, another for emergency responders, and a general-purpose model for all others. The ML logic won’t know which agents are present in the scene (and hence which objects will be needed) until runtime. Caches are too small to hold a copy of every object that might be needed, yet even a single cache miss will leave the task waiting while data is fetched. We are not aware of any SP or ML serving infrastructure with a mix of features matched to these requirements.
To assess the importance of this effect, we ran latency-sensitive ML models on both a local cluster of servers and a public cloud environment using scalable SP server-pool technologies. In all cases we observed frequent pipeline stalls stemming from data accesses that required fetching objects over the datacenter network. These stalls contribute to request backlogs that increase latency, degrade throughput and utilization, and inflate costs.
The correlation grouping mechanism proposed here has two main goals: (i) offer runtime mechanisms whereby developers can express application-specific knowledge about data/computation correlations in a deployment-agnostic way; and (ii) leverage this data to significantly reduce latency, particularly as workload increases and the platform scales out. The mechanism consists of three elements: developer-specified logic that will attach a string, called an correlation key, to each new incoming request; developer-specific code to tag stored data objects with correlation keys; and a runtime engine that makes use of correlation keys to optimize object storage and caching decisions. Leveraging correlation groups does require changes to the runtime platform, but involves minimal changes to the ML application. Benefits include enabling proactive collocation of data with computation, coordinated prefetching, and more efficient cache management.
The remainder of this work is organized as follows. Section 2 motivate our proposals by introducing a representative latency-sensitive ML-based application and identifying shortcomings of modern platforms in face of the application data access patterns. This application is used as a running example throughout this paper. We then propose the correlation grouping mechanism in Section 3. Experimental results showing the benefits of the proposed correlation grouping mechanism are reported in Section 4. Next, we support our motivating claims regarding the shortcomings of modern platforms with experimental results reported in Section 5. Related work is presented in Section 6, and we conclude the paper in Section 7.
2. Motivation
To motivate our effort we first introduce a Real-Time Collision Prediction (RCP) application which is used as a running example in the remainder of the paper. Although we cast it in a somewhat fanciful setting, the workflow embodied in the RCP is representative of a large, important class of latency-sensitive applications (Crankshaw et al., 2020) seen in settings such as factories, robotic warehouses and commercial retail. We then identify limitations of modern SP and ML serving platforms in face of the data access patterns showcased by the RCP application. Finally, we identify trade-offs when addressing these limitations in modern platforms.
2.1. Real-Time Collision Prediction Application
The goal of the application is to continuously predict potential imminent collisions between agents (e.g. cars, pedestrians, cyclists) in traffic intersections with a window of a few seconds. As seen in Figure 1 the RCP pipeline has three steps: multi-object tracking (MOT), trajectory prediction (PRED), and collision detection (CD). The RCP pipeline focuses on the ML, but would be deployed near cameras, signaling danger by flashing large red lights or notifying the hazard warning systems of suitably equipped vehicles. Implementation details are provided in Section 4.1.
Figure 1. Computational graph of the RCP application.
Computational graph of the RCP application.
Figure 1. Computational graph of the RCP application.
The first step, MOT, is responsible for detecting and tracking agents in video frames received from cameras. The computation in this step consists of detecting all agents in each frame and their positions, and then matching each detected agent with its prior instance(s) if any, to determine speed and trajectory. As input, the MOT requires the current frame and the positions and features of agents found in the prior frame. Output is a set of new positions and features for all agents detected in the frame.
A separate instance of the second step, PRED, is triggered for each new position of each agent detected by the MOT step. In this step, the next 12 positions of an agent are predicted based on the past 8 positions of that agent. Thus this step requires the current position of the agent and the previous 7 positions tracked by MOT. The output is thus a trajectory composed of 12 positions.
Each trajectory predicted by PRED triggers a separate instance of the third step, CD. This step matches the received trajectory with all other trajectories predicted so far for the same frame, in order to detect collisions. In addition to the received trajectory, this step also access all available trajectories from the same frame. For each frame, after all instances of CD have been processed, the predicted trajectories of every pair of detected agents will have been evaluated. The output of each instance of this step is a list of collisions.
2.2. Data Access Patterns and Network Overheads
Each step of the RCP application has a different data access pattern. In the MOT step, each request requires a fresh and potentially big data object – features and positions of agents in the previous frame (up to 10MB in our experiments). The nature of the input (a live stream) implies that these objects will only be used once, hence caching would be ineffective. One way to minimize overhead would be to send the object to the node where the next request will be processed. However, the MOT step only learns which objects it requires at runtime, after receiving the request.
PRED requires small but very fresh objects. Although agent positions can be re-used up to 7 times, there are many PRED instances for each frame, thus the overhead from cache misses can add up and become non-negligible in a time-pressured scenario. This overhead grows as more nodes are added to scale out a deployment and cache misses are more common. CD follows a similar rationale: the same trajectory is accessed many times in a short time (the same frame).
We show how significant these overheads can be in the experiments reported in Sections 4 and 5, while more details on modern platforms limitations can be found in Section 6.
2.3. Trade-offs and Limitations
It may seem that we have expressed an impossible objective, but in fact this is not the case. Developers do understand the data dependencies of their ML models, and we will show that they can express them in an simple way. Our central insight is that this is possible, and that the platform can then use the knowledge gained to improve collocation of data and computation. Existing platforms (such as standard cloud frameworks) lack the hooks required to support this capability, in part because of a presumption that such a mechanism would be in tension with other required features such as load balancing and auto-scaling. We will show that with very modest changes to the runtime infrastructure, significant latency improvements are possible even as these dynamic adaptation mechanisms do their job.
To support these claims, we deployed the RCP application twice. We first focused on a platform called Cascade, which is open source and easily extensible. Here the benefits of correlation grouping were dramatic. But then we deployed the identical RCP logic on Microsoft Azure, and ran experiments carefully following recommendations, and using Azure nodes configured for a high quality of service. Azure has fewer extension hooks than Cascade, which prevents us from implementing correlation grouping. Yet following a similar approach yields significant benefit, as shown in Section 5.
3. Correlation Grouping
We now propose the correlation grouping mechanism. As noted in the introduction, the approach has two broad objectives: (i) to offer runtime mechanisms whereby developers can express application-specific knowledge about data/computation correlations in a deployment-agnostic way; and (ii) to leverage this data to significantly reduce latency, particularly as workload increases and the platform scales out. In this section we first discuss requirements for the mechanism that arise from the limitations and trade-offs identified in Section 2. Then we introduce the mechanism per se.
3.1. Requirements
Ideally, developer-oriented mechanisms should be deployment-agnostic: its behavior should be predictable without deployment details. For example, although Apache Spark Resilient Distributed Datasets (RDDs) (Zaharia et al., 2012) offer a way to specify a list of location preferences when storing data, these are treated as hints. The only way to be certain that computation will be collocated with data is to hand-code a manual routing mechanism, using internal APIs to determine which servers have copies of which data items, their current load and request queue length, and resource availability.
Mechanisms such as location preferences in RDD, the partition key in Cosmos DB (Reagan, 2018), and the hash tag in Redis (Redis, 2023) all facilitate collocation of correlated data. For example, data about positions of agents in the same image could be collocated in our RCP application. The cloud runtime environment then introduces further mechanisms intended to collocate  stages of a pipelined computation (for example, inference MOT requests associated with the same RCP video stream) so that a single node will handle them. Other examples of collocating computation include Apache Spark Streaming partitions and Apache Storm stream grouping (Storm, 2023). The problem here is that the data collocation features are decoupled from the computation collocation options, and each service uses its own way of expressing correlations. Developers need a unified mechanism that spans the full deployment and covers both storage placement choices and computational placement decisions.
A mechanism should also be expressive enough to enable the platform to track the data and computation correlations seen at runtime. For example, a single computation may require several data objects, and a single data object may be required by multiple computations. Keeping track of such correlations enables the platform to make optimized decisions regarding data and computation placement. Finally, the mechanism should be lightweight, i.e. the overhead introduced by the mechanism should be negligible and remain constant as the platform scales out.
3.2. Correlation Grouping Mechanism
Our goals and requirements create a complex problem, particularly in modern cloud architectures where one often finds completely independent subsystems specialized for different roles, each with its own design. In particular, storage and computation are generally treated as distinct subsystems, despite the fact that many compute nodes have substantial caches that can be viewed as a component of the storage layer. This observation became the launch point for our correlation grouping mechanism, which groups correlated objects (and hence is a storage-layer feature), yet exposes the locations of these groups so that runtime schedulers can leverage object location data in their decisions.
The next question is how correlation grouping should be expressed. A common approach that we rejected as inflexible assumes that when coding an application, developers can already anticipate the data objects required by each computational stage and declare these dependencies (e.g. using a declarative language or a SQL query) (Zhou et al., 2022a; Makrynioti and Vassalos, 2021; Shaowang et al., 2021). The platform then can parse the specification and create its own internal representation to keep track of which data objects are required by which computations. However, as seen in the RCP application, an ML model might dynamically fetch large objects based on analysis of its inputs. Such model would have an unspecifiable yet latency-critical data dependency.
The alternative approach favored here is to add a label to every request, similar in spirit to hash tags in Redis or partition keys in Azure EHs. The benefit is that labels can be computed dynamically as inputs are classified, so that (for example) an input image showing a taxi could be given a key matching various category-specific ML models and data previously saved in the platform. We call such labels correlation keys: each object has both a unique name (for example, a file pathname or a K/V key) and an correlation key, which would not be unique. Such an approach documents correlations between data and computation: any pair of data/computation with the same correlation key will be understood to be correlated.
The challenge to this approach is that we need to ensure that the mapping between objects and correlation keys can easily be tracked and leveraged across the whole infrastructure. This has to be done in a manner that is lightweight and that can scale to deployments with very large numbers of applications, each with high event rate workflows, and each potentially having a unique correlation key mapping.
Although it can be implemented in different ways, the proposed correlation grouping mechanism affects two aspects of a platform: the application-level API, and the platform-level runtime engine. In the proposed mechanism, the application-level API provides developers with a way to encode application-specific knowledge, while the platform-level runtime engine makes use of the knowledge provided by developers to improve data/computation collocation, scheduling, prefetching, among other benefits discussed in Section 3.3.
The core of the correlation grouping mechanism is a function $f(r)$, which maps any request $r$ to a set of correlation keys. We abstract different operations in a platform as requests: storing/retrieving a data object is a data request and requesting a computational task to be started is a computation request. Correlation keys keys are labels usually implemented as strings. The application-level API allows developers to provide the correlation grouping function $f$ to the platform. The platform then can apply function $f$ to any request: $f$ extracts information from the request to generate a list of one or more keys. Application-specific knowledge is thus entirely encapsulated in $f$. Note that $f$ will be available throughout the distributed service, and must return the same result for a given request no matter where it is invoked. Moreover, it is often invoked on critical paths, where blocking would be problematic. For example, in Section 4 we discuss an correlation grouping function that uses regular expressions on object keys.
In the RCP application, for example, $f$ can map data requests of past positions of an agent $a$ to a label “agent_a”. Computation requests that start the PRED step to predict the future trajectory of agent $a$ can also be mapped to the same label, letting the platform know at runtime that the corresponding objects and the computation are correlated.
The platform-level runtime engine consists of a set of components distributed across the platform that makes use of the developer-provided function $f$. The platform must guarantee at least that: (i) the location where data objects are stored and/or cached depends on the correlation keys; and (ii) the location where computational tasks are placed depends on the correlation keys. We describe next how correlation grouping satisfies the requirements discussed in Section 3.1.
The proposed mechanism is deployment-agnostic, since function $f$ maps requests to correlation keys according only to application-specific knowledge. The platform is then responsible to decide how to handle requests from each label according to the deployment. Since function $f$ applies to both data and computation requests, the platform-level runtime engine is able to consider both data and computation placement in tandem. Computations and data objects are linked through labels, instead of an explicit static list of input objects or query. Developers have a great deal of flexibility when expressing data access patterns of their application.
Turning to the lightweight requirement, we argue that mechanisms such as hash tags and partition keys are potentially costly: the mapping between object keys and labels must be synchronized across the whole infrastructure, which can put a costly distributed operation such as a database update on the critical path. With correlation grouping, only function $f$ itself must be available on all components of a platform: there is no associated replicated state.
3.3. Potential Benefits
A platform-level runtime engine can take advantage of correlation keys in many different ways.
Proactive collocation: It is possible to proactively collocate correlated data and computation by factoring correlation keys into the caching infrastructure, as well as into placement decisions of both computations on servers and objects within a scalable storage infrastructure. For example, in the RCP application, all positions from the same agent (output from the MOT step) can be stored in the same physical node where the PRED computation for that agent will take place when/if it arrives. In this example, the correlation key for all data requests and for the computation request should be the same (the identifier of the agent, for example). Collocation avoids the extra overhead of fetching input data from remote processes. Section 4 shows evidence of this benefit: as seen in Figure 5, end-to-end latency is significantly lower and more consistent when employing correlation grouping.
Prefetching: The knowledge encoded into correlation keys can also enable proactive decisions in anticipation of a future need. For example, in a multistage job a scheduler launching the first stage of a computation might also dispatch a notification that some collection of objects will be required in the second or subsequent stages. This is why our correlation key mechanism permits one object to be tagged with multiple keys: if we consider an ML computation such as RCP with a graphical structure, the output of one stage might be an input to multiple downstream stages, each with its own group of data dependencies. Using correlation keys, this would enable the platform to prefetch objects that will be needed but are not currently in cache, to transfer objects from originating node to a consuming node in anticipation of the future need, etc. Notice the similarity here to a publish-subscribe model, in which publishers might not know who the subscribers will be, and subscribers might not know their publishers. In our work, the correlation key is playing an analogous role.
Consistency: Many edge ML applications are sensitive to data sequencing. Similarly to partition keys in Azure EH, correlation keys can be used to guarantee the consistency of requests. Requests with the same correlation key may have to be handled sequentially and in order. In platforms where objects sharing the same correlation key are to be stored on the same “home nodes”, one update can be used to update the entire set as a single atomic action.
Parallelism: Requests with different correlation keys have no dependency with each other and may thus be handled in parallel. Furthermore, the correlation grouping mechanism enables a finer grain of parallelism control compared, for example, with a mechanism such as Azure EHs. In EHs, requests with different partition keys may be placed in a same partition, causing the requests to be handled sequentially, although they could be potentially have been parallelized.
4. Scaling with Correlation Grouping
In this section, we describe an implementation of the correlation grouping mechanism in Cascade (Song et al., 2022; Derecho Project, 2023): a state-of-the-art stream processing platform targeted at latency-sensitive ML workflows. Cascade was selected because it offered the highest baseline performance among open source platforms amenable to our methodology, enabling us to ask whether even lower latency and higher throughput might be feasible using an correlation grouping methodology.
The goal of this section is to show the benefit of correlation grouping on the end-to-end (E2E) latency of the RCP application. We evaluate the E2E latency of the application with different workloads, while also scaling out the Cascade deployment. Section 5 provides a deeper analysis on the overheads in each step of the pipeline.
4.1. RCP Application Implementation
Our experiment used the Stanford Drone Dataset (SDD) (Robicquet et al., 2016), which contains 60 aerial videos showing campus intersections. The videos include several types of agents (pedestrians, cars, cyclists, skateboarders, buses, and golf carts). Data source “clients” stream these videos at 2.5 FPS. All three steps are implemented in Python using the PyTorch library.
To implement the MOT step, a multi-object tracker (Broström, 2022) that employs YOLO5 for agent detection was used. StrongSORT (Du et al., 2022) and OSNet (Zhou et al., 2022b) are used for re-identification and trajectory tracking. We selected three videos from the SDD for training (little3, hyang5, and gates3). Data is uncompressed, hence each frame is approximately 8MB in size. State data containing positions and features of agents detected in a frame (required to re-identify agents in the following frame) ranges from a few KBs to around 10MB depending on the number of agents in the frame. Frames from the same video must be processed sequentially, but frames from different videos can be processed in parallel by multiple instances of the MOT step. Appendix A includes plots showing the number of agents detected in each frame of each video, as well as the size of the MOT state.
The PRED step employs a trajectory prediction model, YNet (Mangalam et al., 2021) pretrained on the SDD data by the authors (Girase, 2022). PRED requires eight consecutive positions for each agent and makes no prediction if fewer than eight are available. Position objects are small: 10s of bytes. Instances of PRED for the same agent must be processed sequentially. Instances triggered for different agents may be executed in parallel. Notice that the PRED workload depends not only on how many clients are streaming frames, but also on how many agents are detected in each frame.
CD consists of a simple algorithm that performs a linear interpolation on the predicted trajectories of identified agents, outputting a warning if any pair crosses. Instances of this step corresponding to the same frame and client must be processed sequentially to ensure that all pairs of trajectories will be matched with each other. Different frames (even if from the same client) may be processed in parallel.
4.2. Cascade: System Architecture
Cascade is a full-stack platform for high-speed stream processing that prioritizes low latency by hosting data and compute, avoiding copying or locking on critical data paths and leveraging acceleration technologies such as RDMA or DPDK. Backend components are implemented in C++, while client APIs are available in multiple languages (e.g. Java and Python). Cascade consists of a set of nodes interconnected in a complete network. Nodes can be clients or servers. Client nodes may issue requests to any server node. Server nodes implement two subsystems: storage and computation. Each subsystem is further described next. Figure 2 shows an overview of the system architecture.
Figure 2. Cascade architecture.
Cascade’s storage subsystem implements a sharded Key/Value (K/V) object store. Server nodes are logically grouped into disjoint shards with a few nodes each. Prior to our work, each object key was hashed to determine the home shard for the K/V pair. Cascade supports multiples levels of persistence, but our work considered only the two in-memory mode: trigger requests (which cause a task to run but leave no data behind) volatile storage requests (which replicate an object that will then be retained in memory by all members of the home shard). Volatile storage is limited to objects that can fit within memory. In the RCP scenario, for example, object sizes ranged from a few hundred bytes to a few MBs.
Cascade’s User-Defined Logic (UDL) framework is responsible for executing computational tasks. The code for each task is supplied by the developer as a container or in a DLL. In either case, a self-initialization method is expected to register one or more upcallable functions, each associated with a key prefix. On each Cascade server, an upcall will occur if that server receives a K/V pair with a matching key. For example, if a task is registered using the prefix “/RCP/taxis”, then an update to a K/V pair with key “/RCP/taxis/1234” would trigger that task. Cascade’s trigger operation will cause a single upcall to occur, on whichever node the put was sent to. Cascade’s volatile put has two possible behaviors, under developer control: it can upcall on every node in the shard that the key mapped to, or it can do the upcall on one randomly selected node within the shard.
Although a single K/V trigger or volatile put will trigger a task, once the computation is running that task can access additional objects using Cascade’s K/V get API. Moreover, any task can issue a trigger put or a volatile put, enabling pipelined tasks in which each stage initiates the next stage.
The potential overheads in this model are thus: (1) a P2P send operation is used to send a trigger put; (2) an atomic multicast is used to send a volatile put; (3) if a task calls get to access an object that is not homed on the node where it is executing, it must either be fetched from cache or over the network. Even with high-speed networking primitives supported by RDMA or DPDK, network transfers of large objects are costly, so Cascade must minimize the frequency of such events. Nodes have limited capacity, hence a further issue arises if the execution framework needs to enlarge the pool of compute resources. Lacking a warm-start mechanism, Cascade will incur the cost of fetching all needed objects.
Cascade supports a form of resource partitioning called an object pool. The basic idea is as follows: rather than treating all Cascade nodes as a single sharded service, the nodes are instead organized into groups. A single server node can belong to multiple pools. Cascade allows the developer to configure each pool with its own properties, such as the shard size to use, and the degree of data persistence the pool will offer. A persistent pool can be pinned to specific storage hardware (SSD, rotating disk, Optane, etc). A volatile pool can be pinned to host memory, GPU memory, etc. Pools are identified by pathname prefixes: if /x/y is the prefix of a volatile pool holding GPU memory objects, /x/y/z could name a tensor residing within that GPU memory.
4.3. Correlation Grouping on Cascade
Most aspects of the Cascade architecture can easily accommodate the proposed correlation grouping mechanism. As noted earlier, data and computation requests in Cascade consist of trigger, put or get operations. The first two are parameterized by the key and value (uninterpreted byte vector) of the object being transmitted/saved, whereas get takes a key and returns the corresponding object. We implemented function $f$ as a regular expression, which is matched against request keys, but with one simplification: our initial implementation limits each request to a single correlation key  [^1] In future work we will support mappings to multiple correlation keys. A single key suffices for the RCP application, but multiple keys are likely to be important in more complex applications where the output of one compute stage is consumed by multiple follow-on stages, each with its own dependency set.. The correlation key is a substring of the request key, composed of the characters that matched the regular expression. We employed the Hyperscan library (Wang et al., 2019) for matching regular expressions, which introduced a negligible overhead to Cascade’s critical path: according to microbenchmarks, matching the regular expressions employed in the RCP application was under 300 microseconds on average.
To implement the application-level API of the correlation grouping mechanism, we extended the Cascade client API. The extension allows developers to register regular expressions with a specified object pool. Figure 3 shows a code excerpt (in Python) with an example of how object pools are created, with and without employing correlation grouping. The only modification made to the Cascade API was the addition of the optional argument correlation_set_regex to the method that creates an object pool (line 12). Any put or get operation assigns an correlation key to the corresponding object by matching the object key against the regular expression registered in the corresponding object pool (line 18).
1capi = ServiceClientAPI() # Cascade client API
2subgroup_type = "VolatileCascadeStoreWithStringKey"
3subgroup_index = 0
4
5# creating object pool without correlation grouping
6capi.create_object_pool("/no_grouping",
7                        subgroup_type,subgroup_index)
8
9# creating object pool with correlation grouping
10capi.create_object_pool("/grouping",
11                        subgroup_type,subgroup_index,
12                        correlation_set_regex="_[0-9]+")
13
14# putting object in the object pool that is not grouped
15capi.put("/no_grouping/example_1",None)
16
17# putting object in the object pool that is grouped
18capi.put("/grouping/example_1",None) # correlation key is ’_1’
Figure 3. Code excerpt showing how object pools with/without grouping are created and how objects are put/get.
With respect to the platform-level runtime engine, we made two modifications to Cascade: (i) the mapping from key to shard within an object pool is based on the correlation key instead of the request key; and (ii) the correlation key functions and their matching expressions are registered in all nodes. As described above, Cascade previously selected the shard that will handle a request by hashing the object key supplied with the object describing the request. Our modified policy selects the shard by hashing the correlation key. Jointly, these changes ensure that all objects with the same correlation key will be stored (and replicated) in the same shard, and route requests using that same correlation key to this shard.
An open question concerns whether (and how) schedulers should be extended to factor in correlation grouping considerations. Cascade is already designed to send computations to the home nodes of needed objects, or to nodes where those objects are cached: an ideal policy from our perspective.
4.4. Computing Environment
We now describe a series of experiments conducted on our local cluster. Servers in the cluster are equipped with Mellanox ConnectX-4 VPI NIC cards connected to a Mellanox SB7700 InfiniBand switch. This configuration results in a RDMA-capable 100Gbps network backbone. All servers have their clocks synchronized using PTP (Eidson et al., 2002), making timestamps from different servers comparable with sub-millisecond precision. Servers have two configurations, denoted A and B in this work. Configuration A consists of two Intel Xeon Gold 6242 CPUs, 192 GB of memory, and an NVIDIA Tesla T4 GPU. This configuration is commonplace in Edge Cloud environments, where applications such as RCP are often deployed. Configuration B consists of two Intel Xeon E2690 v0 CPUs and 96 GB of memory, but no GPU. All servers run Linux kernel 5.15.0. Cascade nodes were deployed on Docker (version 24.0.6) containers hosted on the cluster servers. Containers were based on Ubuntu 22.04, with CUDA version 11.7.
Cascade was deployed with three object pools, one per step. Cascade nodes in pools responsible for MOT and PRED were deployed on containers hosted on servers with configuration A, since these steps require GPUs for ML inference. Clients and other Cascade nodes were deployed on containers hosted on servers with configuration B. In this section, we report results with a varying number of shards in each pool, as well as a varying number of nodes in each shard. Each configuration of shards/nodes is called a layout.
Although Cascade includes a scheduler, we configured the system to place objects and trigger computations purely by correlation key hashing. This enables us to focus on the degree to which a purely correlation grouping placement of data and computation can improve latency.
4.5. RCP Application Deployment on Cascade
Each step of the RCP application was deployed as a DLL on Cascade’s UDL framework. UDLs are loaded by each Cascade node on startup. ML models are also loaded on startup by the UDLs – all nodes have local access to models and hyperparameters. All required objects (MOT state data, agent positions and trajectory predictions) are stored/retrieved using Cascade K/V store.
Each client is responsible for a different video, and sends complete video frames at a rate of 2.5 FPS. To send a frame, a client puts an object in the K/V store with a key that will trigger a MOT task. The MOT task retrieves the state data from the previous frame, and puts a separate object for each agent position found in the frame. Each agent position put by the MOT task triggers a PRED computation. Each PRED task first retrieves the past positions of the corresponding agent from the K/V store, and puts an object with the trajectory prediction. Each trajectory prediction triggers a CD computation, which first retrieves all trajectory predictions for the same frame available in the K/V store, and then performs the computation. The result of the collision detection is put in the K/V store. The E2E latency is the time elapsed between the client sending the frame, until the result of the last collision detection for that frame is put in the K/V store. All tasks cache in memory all objects they retrieve or create.
We employed two different workloads: single clients (little3, hyang5, and gates3), and three simultaneous clients (little3 + hyang5 + gates3). Each client send 700 frames, and we discard measurements from the first 100 frames. Each experiment is repeated three times. Cascade was completely cleared of objects and object pools before each run of each experiment, and a setup script was executed to create the necessary object pools.
We grouped requests from each step using the correlation grouping mechanism in a different way. For MOT, grouping was based on the identifier of the client, so that all frames from a same client always went to the same shard. For PRED, grouping was based on agent identifiers, and for CD it was based on both the client identifier and the frame number. Table 1 shows all the employed object pools, examples of object keys, which computations are triggered when objects are put in the corresponding object pool (if any), the regular expressions used to group requests (if any), and the resulting correlation keys from matching the regular expressions against the example keys.
Table 1. Object pools, example keys, triggered steps, regular expressions, and correlation keys.
Object Pool Example Key Step Regex Correlation Key
/collision/tracking/cameras /collision/tracking/cameras/little3_42 MOT /[a-zA-Z0-9]+_ /little3_
/collision/tracking/states /collision/tracking/states/little3_42 /[a-zA-Z0-9]+_ /little3_
/collision/tracking/agent_position /collision/tracking/agent_position/little3_7_42 PRED /[a-zA-Z0-9]+_[0-9]+_ /little3_7_
/collision/prediction/agent_prediction /collision/prediction/agent_prediction/little3_42_7 CD /[a-zA-Z0-9]+_[0-9]+_ /little3_42_
/collision/detection /collision/detection/little3_42_7_5
Results reported in this section compare two different placement strategies, random placement and correlation grouping. In the random placement strategy, the RCP application is executed without using correlation grouping, i.e. objects and computations are placed randomly by hashing requests keys – the standard Cascade behavior. In the correlation grouping strategy, we group requests as described above.
4.6. Results
We now report results for experiments with just one client. We varied the number of shards for MOT (1, 3), PRED (1, 3, 5) and CD (1, 3, 5). The number of nodes in each shard was 1. Figure 4 shows a box plot of the E2E latency (in milliseconds) for different layouts and placement strategies. We show only results for client gates3 for the sake of saving space. Results for other clients followed the same pattern. It is possible to see that correlation grouping significantly decreased both median and 75th percentile latencies, except for layout 1/1/1, in which the grouping mechanism had no effect since there was only 1 shard per step. Furthermore, increasing the number of shards did not improve latency with the random placement strategy. In fact, both the median and percentile latencies slightly increased as more shards were added: fetching overhead increased, due to more cache misses.
Figure 4. E2E latency for gates3 on Cascade.
We now increase the workload to three simultaneous clients. Figure 5 shows the E2E latency (in milliseconds) for three simultaneous clients (little3 + hyang5 + gates3), employing different layouts and placement strategies. Latency was significantly lower and more consistent as the deployment scaled out when employing correlation grouping.
Figure 5. E2E latency for three clients on Cascade.
E2E latency for three clients on Cascade.
Figure 5. E2E latency for three clients on Cascade.
In order to further highlight the benefit of collocation, we disabled the object caching performed by each step of our application. Without caching, steps always have to fetch objects from the K/V store. In this context, the main difference between the random placement and correlation grouping strategies is that correlation grouping guarantees that objects being fetched are always stored locally, i.e. in the same Cascade node (and shard) where the computation is running. In the case of random placement, there is no guarantee: objects are placed randomly across shards.
Figure 6 shows the results comparing the two placement strategies with and without the application-level caching, for a workload of three simultaneous clients, and 3/5/5 shards (MOT/PRED/CD steps, respectively), one node per shard. Due to the zero-copy design of Cascade, E2E latency was the same with or without caching when employing correlation grouping, since memory copies are minimized and there is no serialization overhead when an UDL makes a get request to the same node where it is running: keeping objects in the application memory or fetching them locally from the K/V store incurs virtually the same cost. However, in the case of random placement, disabling caching significantly increased latency and reduced throughput. The median E2E latency was off-scale at more than 58 seconds, hence the bar for this case is replaced with an arrow. Throughput was on average 6.7 FPS, thus the pipeline was not able to handle the 7.5 FPS sent by clients collectively.
Figure 6. E2E latency with/without caching on Cascade.
We also evaluated the impact of replication on the E2E latency of the application. Different layouts with more than one node per shard were employed for this evaluation. On Cascade, when a shard has more than one node, any object stored in that shard is replicated to all nodes in the shard, before any computation is triggered. As a result, any computation triggered in such a shard will have local access to the objects in that shard, even if the previous computation (e.g. state of the previous frame) was performed in another node. This behavior can be seen as a form of prefetching, since nodes store objects that will be used in the future before the corresponding request arrives. However it incurs extra latency since the corresponding computation is only triggered after the replication is completed, while a more sophisticated prefetching feature would not prevent the computation to start while prefetching is done in other nodes.
Figure 7 shows results for different layouts. The first group of bars is presented as reference, and show the E2E latency for 3/5/5 shards, with one node per shard (no replication). The next two bars present results for 1/1/1 shards, with different number of nodes per shard (3 and 3/5/5), so there is no difference between random placement and correlation grouping. The last layout has 1/3/3 shards, with 2 nodes per shard, configuring a compromise between having many shards with just one node each and a single shard with many nodes. Results show that replication reduced latency compared to the baseline (first bar). However, employing correlation grouping and multiple shards still results in better E2E latency. We argue that these results show the potential of further investigating the integration of correlation grouping with prefetching and scheduling.
Figure 7. E2E latency with replication on Cascade.
4.7. Insights
Correlation grouping significantly reduces the latency of the RCP application on Cascade, compared to other placement options Cascade is able to support. Additional benefits include better scalability than with random placement, where adding nodes sometimes degraded performance. Performance was steady across varied workloads, suggesting that correlation grouping improves hardware utilization. Changes to the application were primarily to register regular expressions (Table 1 during initialization, as illustrated in Figure 3). Support for correlation grouping required minor changes to Cascade itself, yet enabled the system to use correlation keys to place objects and route requests. Our results on replication indicate that there is potential to improve latency even further by integrating correlation grouping with prefetching.
5. RCP Deployment on a Public Cloud
A natural question to ask is whether our claim that existing SP and ML serving platforms are inadequate is correct. To address this, we now describe a second deployment of RCP on Microsoft Azure Cloud. Granted, Azure is a public cloud whereas Cascade primarily targets a private cluster, but we believe this is a fair comparison because it genuinely represents a state-of-the-art alternative.
The goal of the section is to show evidence that: (i) data access patterns of applications such as RCP indeed result in extra fetching overheads; and (ii) although it is possible to overcome these overheads, it comes at the cost of a high coupling between application and deployment, since there is no unified mechanism available in modern platforms such as the proposed correlation grouping.
5.1. Deployment on Microsoft Azure Cloud
The RCP pipeline was implemented on Stream Analytics (SA) – Azure’s SP platform. We reused the source-code of each of the 3 steps (MOT, PRED, and CD), with few modifications. Each of the ML models was first configured on Azure Machine Learning (AML) as real-time endpoints, using a web-services interface. AML is elastic: depending on load, endpoints can be backed by worker instance pools of varying size (Microsoft, 2022a). Pipeline stages were connected using Azure Event Hubs (EHs) (Kettner and Geisler, 2022). Clients were deployed in a virtual machine. Frames from all three videos were stored in Azure Blob Storage (Calder et al., 2011), which was also used to store MOT state data. Positions of agents and trajectory predictions are stored using Cosmos DB (Reagan, 2018). The resulting architecture is shown in Figure 8, and is compliant with recommendations from the cloud vendor. The individual ML tasks were properly configured with ample resources (similar to the resources employed previously in the Cascade experiments).
Figure 8. RCP deployment on Azure.
Clients stream video frames to the first SA job in the pipeline at a rate of 2.5 FPS, through an EH. The SA job invokes an MOT endpoint: one of the instances in the endpoint is selected by a load balancer to serve the request. The instance fetches the video frame from Blob storage and performs the inference (we assume the client has already uploaded the frame). The state from the previous frame is also fetched from Blob storage if it is not already cached in the instance’s memory. Results are sent through another EH to the next SA job responsible for invoking PRED endpoints for agent positions. For each invocation, the selected instance of the invoked PRED endpoint stores the new position on Cosmos DB, and also fetches the past positions of the corresponding agent that are not already cached in memory. The CD step is performed in a similar fashion. The final output of the pipeline is sent back to the client running on the virtual machine. End-to-end latency of a frame is the time elapsed from when the frame request was first sent by the client, until the last output from the CD step for that frame was received by the client.
All instances in AML endpoints maintain MOT states, positions and trajectory predictions cached in memory for later reuse, thus avoiding fetching from Blob storage or Cosmos DB as much as possible. In the experiments reported in this section, we employ, for each step in the pipeline, a varying number of endpoints and instances per endpoint, as indicated. Instances in MOT and PRED endpoints were of type Standard_NC4as_T4_v3 (equipped with NVIDIA T4 GPUs), while CD instances were of type Standard_DS3_v2.
We employed as many partitions in EHs and streaming units in SA as supported by the endpoints (2*number of instances), maximizing parallelism (Microsoft, 2022b). The partition key for MOT requests was the video name, ensuring that frames originating in any single client were processed sequentially. Similarly, the partition key for PRED was the agent identifier, and for CD the frame number. We set the batch size when invoking AML endpoints to 1, a configuration intended to minimize per-frame latency.
Our experiments stream 700 frames, but we discard measurements the first 100 to give the framework time to warmup. Each experiment was repeated three times. To vary the workload (the number of simultaneous clients streams), we first tested with a single video at a time (little3, hyang5, and gates3), then with two concurrent video streams (little3 + hyang5), and finally with three (little3 + hyang5 + gates3).
5.2. Results: From One to Two Simultaneous Clients
We now report results for experiments with just one client, a case that establishes baseline latency for each video stream, processed individually in a heavily provisioned, dedicated infrastructure. In this set of experiments, each step of the application had one corresponding endpoint, and we varied the number of instances. The MOT endpoint had only one instance, while PRED and CD endpoints had 1, 3, and 5.
Figure 9 shows a box plot of the E2E latencies (in seconds) for each video and number of PRED/CD instances. The number of instances for PRED and CD endpoints is indicated on the horizontal axis. For gates3, the latency for 1 instance was significantly high and thus the bar was replaced by an arrow. The plot shows that increasing the number of instances from 1 to 3 reduced the median and 75th percentiles. However, no significant benefit was observed when increasing the number of instances to 5 for little3 and hyang5. The reason for this is that 3 instances were enough to significantly parallelize invocations to PRED. Increasing the number of instances even further resulted in more network overhead of fetching agent positions that offset the gain of extra parallelism, since the rate of cache misses increased. With 3 instances, 64ms was spent per frame fetching data in the PRED step for hyang5, while 74ms was spent on average with 5 instances.
Figure 9. End-to-end latency for individual clients.
We increased the workload to two simultaneous clients (little3 + hyang5), using the same configuration as the experiments above for individual clients. The average E2E latency, for 5 instances of PRED/CD, was about 67 seconds. The main reason for this significant increase in latency is that a single instance for the MOT step is inadequate to handle the increased rate of incoming frames (5 FPS in the aggregate from the two clients). It is necessary to increase the number of instances of the MOT endpoint.
We conducted experiments with two simultaneous clients, employing 5 instances for PRED and CD endpoints and a varying number of instances for the MOT endpoint (3, 5, 7, and 9). For 3 MOT instances, the E2E latency was significantly high. With 5 to 9 instances, the median latencies were all above 4 seconds, with similar distributions. The benefit of increasing the number of instances was limited, mainly due to the extra network overhead of fetching the MOT states. The extra delay causes requests to pile up in queues, resulting in a significantly higher E2E latency. Figure 10 shows a breakdown of the MOT step for each of the videos, when employing 3 instances on the MOT endpoint. The breakdown shows how much time was spent on average (in ms) on different MOT sub-steps: fetching the video frame, retrieving the state from the previous frame, inference, and uploading the state of the current frame.
Figure 10. Latency breakdown of the MOT step.
5.3. Results: Slashing Overheads From MOT
Since we know that the MOT step requires state data regarding frames from the same video, we deployed a separate MOT endpoint for each video, with one instance each. A separate SA job was also deployed for each endpoint, receiving requests from a separate EH. Then, each client sends requests to the corresponding EH. This ensures that frames from the same video will always be received by the same instance, avoiding the overhead of fetching states since they will be already in memory. We say that MOT step is now grouped. Figure 11 shows the E2E latency of two and three simultaneous clients, employing three MOT endpoints with one instance each, and a varying number of PRED and CD instances (3, 5, and 7). It is possible to see that latency became lower and more consistent as the number of instances for PRED and CD endpoints increased.
Figure 11. Grouped MOT with 2 and 3 simultaneous clients.
However, we observed that there is still potential to further cut network overheads. As we increased the number of PRED and CD instances, parallelism is increased, but so is the time spent fetching agent positions and trajectory predictions from Cosmos DB. Figure 12 shows the latency breakdown for PRED and CD steps, with three simultaneous clients, three grouped MOT endpoints, and a varying number of PRED and CD instances (3, 5, and 7). The plot shows the average time spent per frame (in ms) fetching the necessary input data and running the corresponding inference. The values shown are a sum of the measured times for all invocations. With 7 PRED/CD instances, for example, the average time spent fetching data per frame in PRED and CD steps was almost 200ms (roughly 100ms for each step).
Figure 12. Latency breakdown of PRED and CD steps.
5.4. Results: Slashing Overheads From PRED and CD
Here we employ the same approach as described above for MOT, leveraging application-specific knowledge to group PRED and CD. We deploy multiple endpoints with one instance each, instead of one endpoint with multiple instances. However, it is not possible to rely any longer on the endpoint load balancer to select which instance will receive each request. Thus, we manually select in the SA jobs code which endpoint to forward each request, by writing the output of the previous step to the corresponding EH. For PRED, this selection was based on the agent identifier, modulo the number of PRED endpoints. For CD, the selection was based on the frame number, modulo the number of CD endpoints.
Figure 13 shows a code excerpt (simplified for the sake of presentation) from the SA job responsible for the MOT endpoint associated with the video little3. In the code, from-little3-client is an EH where the client responsible for the little3 video sends requests. Function mot-little3-endpoint invokes the AML endpoint containing a single instance that performs the MOT inference. There are two queries (lines 9 and 12) that forwards the position of each agent to a specific EH (to-pred-0 or to-pred-1) depending on the agent identifier and the number of PRED endpoints. The EHs to-pred-0 and to-pred-1 will deliver the agents positions to the corresponding SA jobs that will invoke their associated PRED endpoints.
1WITH mot_output AS (
2  SELECT [mot-little3-endpoint](input) AS res
3  FROM [from-little3-client] AS input
4), expanded_output AS (
5  SELECT r.arrayvalue.* FROM mot_output
6  CROSS APPLY GetArrayElements(res) AS r
7)
8
9SELECT * INTO [to-pred-0] FROM expanded_output
10WHERE agent_id % num_pred_endpoints = 0
11
12SELECT * INTO [to-pred-1] FROM expanded_output
13WHERE agent_id % num_pred_endpoints = 1
Figure 13. Code excerpt from the MOT SA job for little3.
Figure 14 shows a comparison between the pipeline with only the MOT step grouped and the pipeline with all three steps grouped. We varied the number of endpoints (with one instance each) for PRED and CD (3, 5, and 7, in different combinations), while MOT always had three endpoints (one for each video). As a workload, we configured three side-by-side client streams. Latency for 3 PRED/CD instances and only MOT grouped was significantly high (median 95 seconds) and was replaced by an arrow for the sake of presentation. It is possible to observe a significant benefit when grouping all three steps. The measured latency stabilizes with 5 PRED/CD instances, no benefit was observed when further adding more instances. The gain in parallelism was limited, however, we argue that more instances would be able to handle a heavier workload without an increase in latency.
Figure 14. Grouping only MOT and grouping all steps.
5.5. Insights, Trade-offs, and Limitations
Without our manual grouping, scaling in Azure is straightforward: it is only necessary to increase the number of instances in AML, streaming units in SA, and/or partitions in EHs. AML endpoints will distribute requests across all the available instances automatically according to their load. Furthermore, it is possible to use auto-scaling features, which will scale the deployment in/out according to predefined metrics.
However, when employing the manual grouping strategy, scaling is done by adding or removing endpoints, which requires the application to be reconfigured. In the case of our deployment in Azure, changing endpoints affects client and SA jobs code, and it is also necessary to create new EHs. Load balancing and auto-scaling would need to be implemented by the application itself.
Despite all trade-offs, the benefit of grouping is significant, as also observed in Section 4. For MOT in particular, grouping was essential to support more than one client. Without grouping, the overhead of fetching states caused each MOT request to take more than 400ms, which is the time it takes for each client to send the next frame request. This caused requests to pile up in queues at the first step of the pipeline. In a time-pressured scenario such as the RCP application, the 400ms threshold is of high importance: the ML model employed in the PRED step predicts agents trajectories in the next 4.8 seconds. If it takes too long to process each frame, due to queuing, then the prediction may not be useful any longer. Furthermore, since the application handles a real-time video stream, the MOT state data for each frame will only be ever used once (for the following frame). However, without grouping it is moved twice (upload and download), which can be considered wasteful.
Another clear benefit of grouping is better instance utilization. One could argue that each video/client configures a separate application which could be deployed separately, in a dedicated infrastructure. However, we showed that grouping allows multiple clients to share the same infrastructure while maintaining the same level of performance, which requires less instances to be deployed since they are better utilized. Furthermore, such shared deployment enables new clients to be dynamically added in runtime. However, this would require an extra front-end step in the pipeline that decides how the grouping of a new client should be performed (we’ve done it statically for our experiments).
6. Related Work
Prior efforts (Lee et al., 2018; Kraft et al., 2020; Park et al., 2022; Ahmed et al., 2019; Paganelli et al., 2023) argue for a white box approach, in which optimizations are employed in compilation or training time, based on specific features such as topology of neural networks. While such approaches can reduce the computational cost of model serving, our work employs a black box approach in which the application only needs to identify data / compute correlations. Although the authors in (Lee et al., 2018) criticize black box approaches, they focus only on caching, buffering and batching in non-pipelined applications. The correlation grouping mechanism can support all of these capabilities but also enables additional optimizations, such as proactive data/computation collocation and prefetching.
Existing systems offer a number of mechanisms to express application-specific knowledge and/or request grouping; these features can be found in Redis hash tag, Azure Cosmos DB, EHs and SA partition key, RDD location preferences, Apache Spark Streaming partitions, and Apache Storm stream grouping. However, as noted in Section 3, such mechanisms: (i) are highly coupled with deployment; (ii) do not collocate data and computation in a unified fashion; (iii) are not able to express all possible correlations between data and computation; and/or (iv) may have a non-negligible computational cost as the deployment scales out.
The Pheromone (Yu et al., 2023) system offers a novel approach that the authors refer to as a data bucket abstraction, in which the outputs of functions are automatically grouped into buckets. Developers can then arrange for downstream functions to use data buckets as inputs. This enables Pheromone to place functions close to where the data is stored. The data bucket abstraction is similar to correlation grouping, however outputs in data buckets are always volatile (once consumed, they are garbage collected). The correlation grouping approach is compatible with data persistence, and we would argue that this gives greater flexibility. For example, some outputs may be needed between separate invocations of a pipeline (e.g. past positions of agents in the PRED step of the RCP application); our approach covers such cases but the data bucket model does not. Furthermore, the authors do not consider more complex workflows in which multiple functions may require access to the same output in different stages, or in which a function requires data located in multiple buckets.
Several works improve data locality by scheduling tasks where input data (or most of it) is located (Nishihara et al., 2017; Moritz et al., 2018; Sreekanti et al., 2020; Yu et al., 2021; Boutin et al., 2014; Crankshaw et al., 2020; Zhang et al., 2017), or by scheduling tasks that interact with each other in the same location (node, VM, or container) (Mahgoub et al., 2022; Kotni et al., 2021). These solutions follow a reactive approach: data is first placed without taking into account their correlations, and then tasks are placed where “most of input data” is located. Due to the data access patterns of latency-sensitive ML pipelines, a proactive approach, such as enabled by correlation grouping, may be necessary. We claim that our proposed mechanism is complementary to scheduler-based solutions.
7. Conclusion
We proposed an correlation grouping mechanism that enables developers to express application-specific knowledge of data/computation correlations, and then for platforms to use this information to reduce latency and improve efficiency. Our results show that the proposed mechanism is able to maintain significantly lower and more consistent latency, as well as higher node utilization as the application workload increases and the infrastructure scales out.
Acknowledgements.
We are grateful to Professor Luís Rodrigues for the insightful comments on early versions of this paper. Partial support for our work was provided by AFRL/RY under the SWEC program, Microsoft, Nvidia and Siemens.
Appendix A Application Workload
Figure 15 show the workload of each video (little3, hyang5, and gates3), as well as the aggregated workload for two simultaneous videos (little3 + hyang5), and three simultaneous videos (little3 + hyang5 + gates3). Workload is expressed in number of agents detected in each frame. It is possible to see that little3 has the lowest workload across all 700 frames, while gates3 has the highest – during the first half of the videos, there are more agents in gates3 frames than in the aggregated little3 + hyang5 frames.
Figure 15. Number of agents for different workloads.
Figure 16 shows the size of the MOT state (in MBs) for each frame of each of the three videos employed in our experiments.
Figure 16. Size of the MOT state for each frame of each video.
oter>

### v2

@@
-Low-Latency ML Inference by Grouping Correlated Data Objects and Computation
+Keep Your Friends Close: Leveraging Affinity Groups to Accelerate AI Inference Workflows
 Abstract.
-ML inference workflows often require low latency and high throughput, yet we lack good options for addressing this need.
-Techniques that reduce latency in other streaming settings (such as caching and optimization-driven scheduling) are of limited value because ML data dependencies are often very large and can change dramatically depending on the triggering event.
-In this work, we propose a novel correlation grouping mechanism that makes it easier for developers to express application-specific data access correlations, enabling coordinated management of data objects in server clusters hosting streaming inference tasks.
-Experiments based on a latency-sensitive ML-based application confirm the limitations of standard techniques while showing that our solution yields dramatically better performance.
-The proposed mechanism is able to maintain significantly lower and more consistent latency, achieves higher node utilization as workload and scale-out increase, and yet requires only minor changes to the code implementing the application.
-conference: Preprint; November 2023; arXiv
+AI inference workflows are typically structured as a pipeline or graph of AI programs triggered by events.
+As events occur, the AIs perform inference or classification tasks under time pressure to respond or take some action.
+Standard techniques that reduce latency in other streaming settings (such as caching and optimization-driven scheduling) are of limited value because AI data access patterns (models, databases) change depending on the triggering event: a significant departure from traditional streaming.
+In this work, we propose a novel affinity grouping mechanism that makes it easier for developers to express application-specific data access correlations, enabling coordinated management of data objects in server clusters hosting streaming inference tasks.
+Our proposals are thus complementary to other approaches such as caching and scheduling.
+Experiments confirm the limitations of standard techniques, while showing that the proposed mechanism is able to maintain significantly lower latency as workload and scale-out increase, and yet requires only minor code changes.
+journalyear: 2025copyright: acmlicensedconference: The 18th ACM International Systems and Storage Conference; September 8–9, 2025; Virtual, Israelbooktitle: The 18th ACM International Systems and Storage Conference (SYSTOR ’25), September 8–9, 2025, Virtual, Israeldoi: 10.1145/3757347.3759129isbn: 979-8-4007-2119-9/2025/09
 1. Introduction
 Latency-sensitive AI-based applications are increasingly common (Chen et al., 2021).
-For instance, in edge intelligence applications, devices pass captured data through Machine Learning (ML) inference pipelines (Deng et al., 2020).
-Minimizing latency is the highest priority, although high throughput and efficient platform utilization remain important as secondary goals (Crankshaw et al., 2017).
+For instance, in edge intelligence applications, devices pass captured data through AI inference and classification pipelines, frequently combined with continuous learning (Deng et al., 2020).
+Minimizing latency is the priority, although throughput and resource utilization remain important goals (Crankshaw et al., 2017).
 Our work starts with the observation that when these pipelines are deployed onto standard Stream Processing (SP) platforms (Isah et al., 2019), a mismatch occurs.
-SP systems need low latency but are typically optimized for highly parallel stateless computations and data transformations.
-ML inference is stateful and often context-sensitive, resulting in data dependencies that can defeat the assumptions made in SP schedulers (Shaowang et al., 2021).
+SP systems are typically optimized for highly parallel stateless computations and data transformations.
+AI inference and classifiers are often stateful and context-sensitive, resulting in data dependencies that can defeat the assumptions made in SP schedulers (Shaowang et al., 2021).
 The issue is further complicated by unpredictability.
-ML models depend on pretrained parameter vectors that can be large.
-For example, a trajectory-computing task for mobile “agents” in a traffic scene might have one set of models trained for cars, another for bicycles, a third for pedestrians, etc.
+For example, a trajectory-computing task for mobile “actors” in a traffic scene might have different models trained for cars, bicycles, pedestrians, etc.
 Among the vehicle models there could be one specialized for taxis, another for emergency responders, and a general-purpose model for all others.
-The ML logic won’t know which agents are present in the scene (and hence which objects will be needed) until runtime.
-Caches are too small to hold a copy of every object that might be needed, yet even a single cache miss will leave the task waiting while data is fetched.
-We are not aware of any SP or ML serving infrastructure with a mix of features matched to these requirements.
-To assess the importance of this effect, we ran latency-sensitive ML models on both a local cluster of servers and a public cloud environment using scalable SP server-pool technologies.
-In all cases we observed frequent pipeline stalls stemming from data accesses that required fetching objects over the datacenter network.
-These stalls contribute to request backlogs that increase latency, degrade throughput and utilization, and inflate costs.
-The correlation grouping mechanism proposed here has two main goals: (i) offer runtime mechanisms whereby developers can express application-specific knowledge about data/computation correlations in a deployment-agnostic way; and (ii) leverage this data to significantly reduce latency, particularly as workload increases and the platform scales out.
-The mechanism consists of three elements: developer-specified logic that will attach a string, called an correlation key, to each new incoming request; developer-specific code to tag stored data objects with correlation keys; and a runtime engine that makes use of correlation keys to optimize object storage and caching decisions.
-Leveraging correlation groups does require changes to the runtime platform, but involves minimal changes to the ML application.
+The AI won’t know which actors are present in the scene (and hence which objects will be needed) until runtime.
+Caches are too small to hold a copy of every object that might be needed, yet even a single cache miss will leave the task waiting while data is fetched over the network.
+Data movement overheads are further amplified when different stages of an AI pipeline are executed in different nodes of a network.
+Reasons for distributing different stages across multiple nodes include scalability, load balancing, parallelism, and specialized hardware sharing.
+Scalable storage tends to randomly distribute data objects across the infrastructure.
+Although schedulers try to minimize data movement when deciding where to execute a pipeline stage, there is no collocation guarantee.
+AI developers understand well which collocation dependencies would be beneficial for the latency of their applications.
+Alas, existing platforms (such as standard cloud frameworks) lack the hooks required for developers to provide feedback.
+We conjecture that this is partly because of a presumption that such a mechanism would be in tension with core platform features such as load balancing and auto-scaling, and in part because of the concern that the annotation would require significant effort from the developers.
+In this work, we propose the affinity grouping mechanism, which has two main goals: (i) to offer standardized, platform-independent annotation method whereby developers can express application-specific knowledge about data/computation correlations in a deployment-agnostic way (allowing mechanisms such as auto-scaling and load balancing to keep doing their job); and (ii) leverage this data to significantly reduce latency as workload increases and the platform scales out.
+The mechanism consists of three elements: developer-specified logic that attaches a string, called an affinity key, to each new incoming request; developer-specific code to tag stored data objects with affinity keys; and a runtime engine that makes use of affinity keys to optimize object and task placement decisions.
+Leveraging affinity groups does require changes to the runtime platform, but involves minimal changes to the AI application.
 Benefits include enabling proactive collocation of data with computation, coordinated prefetching, and more efficient cache management.
+Our proposals are thus complementary to other approaches such as caching and scheduling.
+In order to evaluate our proposals, we design and implement a representative latency-sensitive AI-based application consisting of a composition of off-the-shelf AI models.
+Using this application, two sets of experiments are performed.
+First, we deploy our representative application on a state-of-the-art stream processing platform targeted at latency-sensitive AI workflows, which is based on a K/V store.
+Results show that affinity grouping achieves significantly lower and more consistent latency as workload and scale-out increase, compared to the standard object and task placement strategy of the employed platform.
+All this requires a minimal annotation effort from the developer.
+In the second set of experiments, we deploy the same application on a public cloud, Microsoft Azure.
+The goal is to show evidence of the data movement overheads present in existing SP and AI serving platforms.
+We observe frequent pipeline stalls stemming from data accesses that required fetching objects over the network.
+We then modify our application to address the observed overheads, at the cost of a high coupling between application and deployment.
+Figure 1. Computational graph of the RCP application.
+[…]
+Computational graph of the RCP application.
+This work makes the following contributions:
+  • We identify inefficiencies of modern platforms to collocate data when deploying a representative latency-sensitive AI-based application that comprises off-the-shelf AI models.
+We support this claim by investigating the application performance on a public cloud.
+  • To address the identified issue, we propose affinity grouping, an easy-to-use mechanism that requires no developer knowledge of deployment/environment details, and yet enables platforms to achieve effective collocation of data and compute.
+  • An implementation of the affinity grouping mechanism on a state-of-the-art AI serving platform based on a K/V store is described and evaluated.
 The remainder of this work is organized as follows.
-Section 2 motivate our proposals by introducing a representative latency-sensitive ML-based application and identifying shortcomings of modern platforms in face of the application data access patterns.
+Section 2 motivate our proposals by introducing a representative latency-sensitive AI-based application and identifying shortcomings of modern platforms in face of the application data access patterns.
 This application is used as a running example throughout this paper.
-We then propose the correlation grouping mechanism in Section 3.
-Experimental results showing the benefits of the proposed correlation grouping mechanism are reported in Section 4.
-Next, we support our motivating claims regarding the shortcomings of modern platforms with experimental results reported in Section 5.
-Related work is presented in Section 6, and we conclude the paper in Section 7.
+We then propose the affinity grouping mechanism in Section 3.
+Experimental results showing the benefits of the proposed affinity grouping mechanism on a local cluster are reported in Section 4.
+A series of experiments on Microsoft Azure that support our motivating claims are reported in Section 5.
+Related work is presented in Section 6, while Section 7 discuss the generality of our approach.
+We conclude the paper in Section 8.
 2. Motivation
-To motivate our effort we first introduce a Real-Time Collision Prediction (RCP) application which is used as a running example in the remainder of the paper.
-Although we cast it in a somewhat fanciful setting, the workflow embodied in the RCP is representative of a large, important class of latency-sensitive applications (Crankshaw et al., 2020) seen in settings such as factories, robotic warehouses and commercial retail.
-We then identify limitations of modern SP and ML serving platforms in face of the data access patterns showcased by the RCP application.
-Finally, we identify trade-offs when addressing these limitations in modern platforms.
-2.1. Real-Time Collision Prediction Application
-The goal of the application is to continuously predict potential imminent collisions between agents (e.g.
-cars, pedestrians, cyclists) in traffic intersections with a window of a few seconds.
+We first introduce a Real-Time Collision Prediction (RCP) application that is used as a running example in this paper.
+RCP is representative of a large, important class of latency-sensitive applications (Crankshaw et al., 2020) seen in settings such as factories, robotic warehouses and commercial retail.
+We then identify limitations of modern SP and AI serving platforms in face of the data access patterns showcased by the RCP application.
+2.1. Real-Time Collision Prediction
+The RCP application continuously sense imminent collisions between actors (e.g.
+cars, pedestrians) in traffic intersections with a window of a few seconds.
+It consists of a composition of off-the-shelf AI models (trained for different tasks), resulting in a pipeline that runs in an edge datacenter, near traffic cameras.
+We focus on the AI inference pipeline, but examples of reactions to the pipeline output could include flashing red lights or notifying suitably equipped vehicles.
 As seen in Figure 1 the RCP pipeline has three steps: multi-object tracking (MOT), trajectory prediction (PRED), and collision detection (CD).
-The RCP pipeline focuses on the ML, but would be deployed near cameras, signaling danger by flashing large red lights or notifying the hazard warning systems of suitably equipped vehicles.
 Implementation details are provided in Section 4.1.
-Figure 1. Computational graph of the RCP application.
-Computational graph of the RCP application.
-Figure 1. Computational graph of the RCP application.
-The first step, MOT, is responsible for detecting and tracking agents in video frames received from cameras.
-The computation in this step consists of detecting all agents in each frame and their positions, and then matching each detected agent with its prior instance(s) if any, to determine speed and trajectory.
-As input, the MOT requires the current frame and the positions and features of agents found in the prior frame.
-Output is a set of new positions and features for all agents detected in the frame.
-A separate instance of the second step, PRED, is triggered for each new position of each agent detected by the MOT step.
-In this step, the next 12 positions of an agent are predicted based on the past 8 positions of that agent.
-Thus this step requires the current position of the agent and the previous 7 positions tracked by MOT.
-The output is thus a trajectory composed of 12 positions.
+The first step, MOT, is responsible for detecting and tracking actors in video frames received from cameras.
+The computation in this step consists of detecting all actors in each frame and their positions, and then matching each detected actor with its prior instance(s) if any, to determine speed and trajectory.
+As input, the MOT requires the current frame and the positions and features of actors found in the prior frame.
+Output is a set of new positions and features for all actors detected in the frame.
+A separate instance of the second step, PRED, is triggered for each new position of each actor detected by MOT.
+PRED predicts the next $q$ positions of an actor based on the past $p$ positions of that actor.
+Thus this step requires the current position of the actor and the previous p1p-1 positions tracked by MOT.
+The output is thus a trajectory composed of $q$ positions.
+In our implementation $p=8$ and $q=12$, since the pretrained model employed was trained with those values.
 Each trajectory predicted by PRED triggers a separate instance of the third step, CD.
@@
 In addition to the received trajectory, this step also access all available trajectories from the same frame.
-For each frame, after all instances of CD have been processed, the predicted trajectories of every pair of detected agents will have been evaluated.
+For each frame, after all instances of CD have been processed, the predicted trajectories of every pair of detected actors will have been evaluated.
 The output of each instance of this step is a list of collisions.
-2.2. Data Access Patterns and Network Overheads
-Each step of the RCP application has a different data access pattern.
-In the MOT step, each request requires a fresh and potentially big data object – features and positions of agents in the previous frame (up to 10MB in our experiments).
-The nature of the input (a live stream) implies that these objects will only be used once, hence caching would be ineffective.
-One way to minimize overhead would be to send the object to the node where the next request will be processed.
+2.2. RCP Distributed Deployment
+Earlier, we noted that applications such as our RCP example lend themselves to a distributed deployment.
+What we failed to say is that on many platforms, doing so would be a big departure from what the platform supports.
+In fact, today’s most common approach for implementing an application like the RCP is to combine the whole pipeline into a monolithic application, then run it on one multi-core server, selected by a load balancer.
+In order to scale and keep up in real-time, load balancing tends to randomly spray requests over nodes.
+Although this approach avoids overheads associated with transitioning from one step to the next (since all steps happen in the same process), with complex tasks such as the RCP, it could easily exceed what one server can handle.
+Consecutive invocations of the pipeline for frames from the same camera have a data dependency.
+In addition, multiple steps (MOT and PRED) require GPU accelerators.
+As a consequence, a frame $k$ must wait until the whole pipeline for frame k1k-1 finishes processing and make the necessary input data available.
+Even frames from other cameras will need GPU resources to be freed.
+Furthermore, multiple instances of PRED, one for each actor detected by MOT in the same frame, can be executed in parallel.
+However, in the monolithic approach they would bottleneck, significantly increasing end-to-end latency for that frame.
+Finally, certain steps may require more resources than others.
+In the monolithic approach, scaling individual steps becomes a challenge.
+For example, in our experiments the PRED step presented the heaviest workload, since each frame had up to 49 actors, and thus this step required more resources.
+Thus a job like RCP lends itself to multi-server parallelism: now the scheduler can place different steps on different servers, taking care not to overload any single server.
+However, latency will still be sensitive to data movement overheads related to the data access patterns of each step.
+We further discuss such overheads next.
+2.3. RCP Data Access Patterns and Overheads
+When we deploy an AI pipeline on multiple servers, load balancing will often run each new task on a server picked at random among lightly loaded machines.
+Yet because AI models and data dependencies can be very large, the new task then pauses to fetch data over the network.
+In real-time applications such as the RCP, standard solutions such as caching are of limited value due to data freshness.
+We discuss next the extent of this effect, and how developer-provided knowledge can help the platform.
+In the MOT step, each request requires a fresh and potentially big data object – features and positions of actors in the previous frame (up to 10MB in our experiments).
+The nature of the input (a live stream) implies that these objects will only be used once, hence caching would be ineffective, since cached objects will never be reused.
+One way to minimize this overhead would be to send the object to the node where the next request will be processed.
 However, the MOT step only learns which objects it requires at runtime, after receiving the request.
 PRED requires small but very fresh objects.
-Although agent positions can be re-used up to 7 times, there are many PRED instances for each frame, thus the overhead from cache misses can add up and become non-negligible in a time-pressured scenario.
-This overhead grows as more nodes are added to scale out a deployment and cache misses are more common.
-CD follows a similar rationale: the same trajectory is accessed many times in a short time (the same frame).
-We show how significant these overheads can be in the experiments reported in Sections 4 and 5, while more details on modern platforms limitations can be found in Section 6.
-2.3. Trade-offs and Limitations
-It may seem that we have expressed an impossible objective, but in fact this is not the case.
-Developers do understand the data dependencies of their ML models, and we will show that they can express them in an simple way.
-Our central insight is that this is possible, and that the platform can then use the knowledge gained to improve collocation of data and computation.
-Existing platforms (such as standard cloud frameworks) lack the hooks required to support this capability, in part because of a presumption that such a mechanism would be in tension with other required features such as load balancing and auto-scaling.
-We will show that with very modest changes to the runtime infrastructure, significant latency improvements are possible even as these dynamic adaptation mechanisms do their job.
-To support these claims, we deployed the RCP application twice.
-We first focused on a platform called Cascade, which is open source and easily extensible.
-Here the benefits of correlation grouping were dramatic.
-But then we deployed the identical RCP logic on Microsoft Azure, and ran experiments carefully following recommendations, and using Azure nodes configured for a high quality of service.
-Azure has fewer extension hooks than Cascade, which prevents us from implementing correlation grouping.
-Yet following a similar approach yields significant benefit, as shown in Section 5.
-3. Correlation Grouping
-We now propose the correlation grouping mechanism.
-As noted in the introduction, the approach has two broad objectives: (i) to offer runtime mechanisms whereby developers can express application-specific knowledge about data/computation correlations in a deployment-agnostic way; and (ii) to leverage this data to significantly reduce latency, particularly as workload increases and the platform scales out.
-In this section we first discuss requirements for the mechanism that arise from the limitations and trade-offs identified in Section 2.
-Then we introduce the mechanism per se.
+Actor positions can be re-used only up to 7 times, and there are many PRED instances for each frame (one for each actor detected).
+Thus the overhead from cache misses can add up and become non-negligible in a time-pressured scenario, as such instances are placed on different servers by the load balancer.
+This overhead grows as more nodes are added to scale out and cache misses are more common.
+Furthermore, which objects will be necessary are only known at runtime, after MOT is finished.
+CD follows a similar rationale: the same trajectory is accessed many times in a short time (the same frame), and the actual workload is not predictable.
+Despite the freshness and the unpredictability of data access, the developer of the RCP application knows the mapping between future requests and objects stored previously.
+A platform can benefit from this knowledge and collocate correlated objects and requests proactively, slashing data movement overheads.
+We achieve that with the affinity grouping mechanism, proposed next.
+Later, we show how significant the overheads discussed above can be in the experiments reported in Sections 4 and 5.
+3. Affinity Grouping
+In this section, we start with an overview of how data and tasks are processed in modern SP and AI serving platforms.
+We then discuss requirements for the mechanism arising from the challenges identified in Section 2.
+Finally, we introduce and discuss the affinity grouping mechanism.
+3.1. Execution and Data Flow Model
+We now describe a general execution and data flow model for a typical platform that hosts applications such as the RCP.
+Our proposals are related to the location at which data is stored/retrieved, and computational tasks are executed.
+A location refers to a network endpoint (e.g.
+a server node) that holds resources such as CPUs, GPUs, hard disks, and memory.
+A platform consists of multiple of such resources distributed across an edge/cloud datacenter network.
+Platforms can be divided in two main subsystems that constitute the platform-level runtime engine: storage and compute.
+As an application runs, data objects are stored and retrieved (e.g.
+actors positions in RCP), and computational tasks are initiated (e.g.
+a step of the RCP pipeline).
+In both cases, the corresponding subsystem needs to make placement decisions.
+The storage subsystem contains a data scheduling component responsible for deciding on the location a data object is to be stored at or retrieved from.
+Similarly, the compute subsystem has a task scheduling component that decides upon the location at which a computational task is started.
+Data objects and/or computational tasks are said to be collocated when the corresponding scheduling components place them at the same location.
+In addition, a platform also provides developers with an application-level API, used to interact with the platform-level runtime engine.
+Garbage collection, i.e.
+how data objects are cleaned from storage when they are not needed (e.g.
+actor positions in RCP are not needed any longer after some seconds), is orthogonal to this work and thus out of scope.
 3.2. Requirements
-Ideally, developer-oriented mechanisms should be deployment-agnostic: its behavior should be predictable without deployment details.
-For example, although Apache Spark Resilient Distributed Datasets (RDDs) (Zaharia et al., 2012) offer a way to specify a list of location preferences when storing data, these are treated as hints.
-The only way to be certain that computation will be collocated with data is to hand-code a manual routing mechanism, using internal APIs to determine which servers have copies of which data items, their current load and request queue length, and resource availability.
-Mechanisms such as location preferences in RDD, the partition key in Cosmos DB (Reagan, 2018), and the hash tag in Redis (Redis, 2023) all facilitate collocation of correlated data.
-For example, data about positions of agents in the same image could be collocated in our RCP application.
-The cloud runtime environment then introduces further mechanisms intended to collocate  stages of a pipelined computation (for example, inference MOT requests associated with the same RCP video stream) so that a single node will handle them.
+Introduction of feedback by the application developer affects two aspects of a platform: the application-level API, and the platform-level runtime engine.
+The API provides developers with a way to encode application-specific knowledge about correlations, while the platform-level runtime engine makes use of the knowledge provided by developers to improve data/computation collocation, scheduling, prefetching, among other benefits discussed in Section 3.4.
+Ideally, the method of specifying correlations through the application-level API exposed by the platform should be deployment-agnostic: the developer logic should not be intertangled with specific deployment details.
+For example, the only way to be certain that a computation in Apache Spark Resilient Distributed Datasets (RDDs) (Zaharia et al., 2012) will be collocated with data is to manually code a routing mechanism, using internal RDDs APIs to determine which servers have copies of which data items.
+A similarly complicated proprietary approach is required to ensure collocation in Microsoft Azure, as we show in Section 5.
+While RDDs offers a way to specify a list of location preferences when storing data, these are only treated as hints, thus not providing a guarantee.
+The method of specifying correlations should also be expressive enough to allow the platform to capture the data and computation correlations at runtime.
+For example, a single computation may require several data objects, and a single data object may be required by multiple computations.
+A common approach that we reject as inflexible assumes that when coding an application, developers can already anticipate the data objects required by each computational stage and statically declare these dependencies (e.g.
+using a declarative language or SQL) (Zhou et al., 2022a; Makrynioti and Vassalos, 2021; Shaowang et al., 2021).
+The platform then can parse the specification and create its own internal representation to keep track of which data objects are required by which computations.
+However, as seen in the RCP application, an AI model might dynamically fetch large objects based on analysis of its inputs.
+Such model would have an unspecifiable yet latency-critical data dependency.
+Mechanisms such as location preferences in RDD (Zaharia et al., 2012), the partition key in Cosmos DB (Reagan, 2018), and the hash tag in Redis (Redis, 2023) all facilitate collocation of correlated data.
+For example, positions of actors in the same image could be collocated in our RCP application.
+The cloud runtime environment then introduces further mechanisms intended to collocate stages of a pipelined computation (for example, MOT inference requests associated with the same RCP video stream) so that a single node handles them.
 Other examples of collocating computation include Apache Spark Streaming partitions and Apache Storm stream grouping (Storm, 2023).
-The problem here is that the data collocation features are decoupled from the computation collocation options, and each service uses its own way of expressing correlations.
+The problem here is that the data collocation features are decoupled from the computation collocation options.
 Developers need a unified mechanism that spans the full deployment and covers both storage placement choices and computational placement decisions.
-A mechanism should also be expressive enough to enable the platform to track the data and computation correlations seen at runtime.
-For example, a single computation may require several data objects, and a single data object may be required by multiple computations.
-Keeping track of such correlations enables the platform to make optimized decisions regarding data and computation placement.
-Finally, the mechanism should be lightweight, i.e.
-the overhead introduced by the mechanism should be negligible and remain constant as the platform scales out.
-3.2. Correlation Grouping Mechanism
+Furthermore, it is important that the specification method is easy-to-use for application developers, i.e.
+only minimal changes to the application code should be necessary.
+Modern APIs for developing AI-based applications generally satisfy this requirement.
+Finally, the mechanism should be lightweight: the engine should support specified collocations efficiently at runtime, i.e.
+the overhead introduced by the implementation within the engine should be negligible and remain constant as the platform scales out.
+Mechanisms such as Redis hash tags and Azure partition keys are potentially costly: the mapping between object keys and labels must be synchronized across the whole infrastructure, which can put a costly distributed operation such as a database update on the critical path.
+Each of the requirements described above is not particularly challenging by itself.
+The challenge comes when trying to satisfy all requirements together.
+For instance, the requirement of expressiveness may be in conflict with ease-of-use: providing more information may also increase the complexity of the API for developers.
+Similarly, designing a unified mechanism may also increase the complexity for developers, and also for the platform to keep track and synchronize correlation information across the infrastructure.
+Moreover, a deployment-agnostic mechanism may not be expressive enough, since deployment details have to be abstracted away.
+3.3. Affinity Grouping Mechanism
 Our goals and requirements create a complex problem, particularly in modern cloud architectures where one often finds completely independent subsystems specialized for different roles, each with its own design.
 In particular, storage and computation are generally treated as distinct subsystems, despite the fact that many compute nodes have substantial caches that can be viewed as a component of the storage layer.
-This observation became the launch point for our correlation grouping mechanism, which groups correlated objects (and hence is a storage-layer feature), yet exposes the locations of these groups so that runtime schedulers can leverage object location data in their decisions.
-The next question is how correlation grouping should be expressed.
-A common approach that we rejected as inflexible assumes that when coding an application, developers can already anticipate the data objects required by each computational stage and declare these dependencies (e.g.
-using a declarative language or a SQL query) (Zhou et al., 2022a; Makrynioti and Vassalos, 2021; Shaowang et al., 2021).
-The platform then can parse the specification and create its own internal representation to keep track of which data objects are required by which computations.
-However, as seen in the RCP application, an ML model might dynamically fetch large objects based on analysis of its inputs.
-Such model would have an unspecifiable yet latency-critical data dependency.
-The alternative approach favored here is to add a label to every request, similar in spirit to hash tags in Redis or partition keys in Azure EHs.
-The benefit is that labels can be computed dynamically as inputs are classified, so that (for example) an input image showing a taxi could be given a key matching various category-specific ML models and data previously saved in the platform.
-We call such labels correlation keys: each object has both a unique name (for example, a file pathname or a K/V key) and an correlation key, which would not be unique.
-Such an approach documents correlations between data and computation: any pair of data/computation with the same correlation key will be understood to be correlated.
-The challenge to this approach is that we need to ensure that the mapping between objects and correlation keys can easily be tracked and leveraged across the whole infrastructure.
-This has to be done in a manner that is lightweight and that can scale to deployments with very large numbers of applications, each with high event rate workflows, and each potentially having a unique correlation key mapping.
-Although it can be implemented in different ways, the proposed correlation grouping mechanism affects two aspects of a platform: the application-level API, and the platform-level runtime engine.
-In the proposed mechanism, the application-level API provides developers with a way to encode application-specific knowledge, while the platform-level runtime engine makes use of the knowledge provided by developers to improve data/computation collocation, scheduling, prefetching, among other benefits discussed in Section 3.3.
-The core of the correlation grouping mechanism is a function $f(r)$, which maps any request $r$ to a set of correlation keys.
-We abstract different operations in a platform as requests: storing/retrieving a data object is a data request and requesting a computational task to be started is a computation request.
-Correlation keys keys are labels usually implemented as strings.
-The application-level API allows developers to provide the correlation grouping function $f$ to the platform.
-The platform then can apply function $f$ to any request: $f$ extracts information from the request to generate a list of one or more keys.
+This observation became the launch point for our affinity grouping mechanism, which groups correlated objects (and hence is a storage-layer feature), yet exposes the locations of these groups so that runtime schedulers can leverage object location data in their decisions.
+In affinity grouping, each data object and computational task has a label, similar in spirit to hash tags in Redis or partition keys in Azure EHs.
+The critical difference is that labels can be computed dynamically at runtime as inputs are classified.
+For example, an input image showing a taxi could be given a label matching various category-specific AI models and data previously saved in the platform.
+We call such labels affinity keys: each object and task has both a unique name (for example, a file pathname or a K/V key) and an affinity key, which would not be unique.
+Such an approach allows the platform to derive correlations between data and computation: any data objects and a task with the same affinity key will be regarded as requiring correlation.
+The core of the proposed mechanism is a function $f(d)$, which maps a descriptor $d$ to an affinity key.
+A descriptor contains metadata about a data object (to be stored or retrieved) or a computational task (to be initiated).
+Affinity keys are labels that can be implemented e.g.
+as strings.
+The application-level API allows developers to provide the affinity grouping function $f$ to the platform.
+When a placement decision for a data object or a computational task needs to be made, the platform then can apply function $f$: the function extracts information from the given descriptor to generate an affinity key.
 Application-specific knowledge is thus entirely encapsulated in $f$.
-Note that $f$ will be available throughout the distributed service, and must return the same result for a given request no matter where it is invoked.
+Note that $f$ will be available throughout the distributed service, and must return the same result for a given descriptor no matter where it is invoked.
 Moreover, it is often invoked on critical paths, where blocking would be problematic.
-For example, in Section 4 we discuss an correlation grouping function that uses regular expressions on object keys.
-In the RCP application, for example, $f$ can map data requests of past positions of an agent $a$ to a label “agent_a”.
-Computation requests that start the PRED step to predict the future trajectory of agent $a$ can also be mapped to the same label, letting the platform know at runtime that the corresponding objects and the computation are correlated.
-The platform-level runtime engine consists of a set of components distributed across the platform that makes use of the developer-provided function $f$.
-The platform must guarantee at least that: (i) the location where data objects are stored and/or cached depends on the correlation keys; and (ii) the location where computational tasks are placed depends on the correlation keys.
-We describe next how correlation grouping satisfies the requirements discussed in Section 3.1.
-The proposed mechanism is deployment-agnostic, since function $f$ maps requests to correlation keys according only to application-specific knowledge.
-The platform is then responsible to decide how to handle requests from each label according to the deployment.
-Since function $f$ applies to both data and computation requests, the platform-level runtime engine is able to consider both data and computation placement in tandem.
-Computations and data objects are linked through labels, instead of an explicit static list of input objects or query.
-Developers have a great deal of flexibility when expressing data access patterns of their application.
-Turning to the lightweight requirement, we argue that mechanisms such as hash tags and partition keys are potentially costly: the mapping between object keys and labels must be synchronized across the whole infrastructure, which can put a costly distributed operation such as a database update on the critical path.
-With correlation grouping, only function $f$ itself must be available on all components of a platform: there is no associated replicated state.
+For example, in Section 4 we discuss an affinity grouping function that uses regular expressions.
+The platform-level runtime engine must guarantee that the location where data objects are stored and/or cached, and computational tasks are placed, depends on affinity keys.
+In the RCP application, for example, $f$ can map data objects of past positions of an actor $a$ to a label “actor_a”.
+The computation task, corresponding to the PRED step, predicting the future trajectory of actor $a$ can also be mapped to the same label, letting the platform know at runtime that the corresponding objects and the computation are correlated.
+Regarding the requirements defined above, the proposed mechanism is deployment-agnostic, since function $f$ maps descriptors to affinity keys according only to application-specific knowledge.
+The platform is then responsible to decide how to handle objects and tasks from each label according to the deployment.
+The mechanism is unified, since function $f$ applies to placement decisions on both storage and compute subsystems.
+Tasks and objects are linked through labels, instead of an explicit static list of input objects or query.
+Developers have a great deal of flexibility when expressing data access patterns of their application (i.e.
+expressive).
+Turning to the efficiency requirement, only function $f$ itself must be available on all components of a platform: there is no associated replicated state.
+Compliance with the ease-of-use requirement, however, depends on the specific implementation.
+Section 4 describes an implementation in which developers only need to provide a regular expression.
 3.4. Potential Benefits
-A platform-level runtime engine can take advantage of correlation keys in many different ways.
-Proactive collocation: It is possible to proactively collocate correlated data and computation by factoring correlation keys into the caching infrastructure, as well as into placement decisions of both computations on servers and objects within a scalable storage infrastructure.
-For example, in the RCP application, all positions from the same agent (output from the MOT step) can be stored in the same physical node where the PRED computation for that agent will take place when/if it arrives.
-In this example, the correlation key for all data requests and for the computation request should be the same (the identifier of the agent, for example).
+Figure 2. Cascade architecture.
+A platform-level runtime engine can take advantage of affinity keys in many different ways.
+Proactive collocation: It is possible to proactively collocate correlated data and computation by factoring affinity keys into the caching infrastructure, as well as into placement decisions of both computations on servers and objects within a scalable storage infrastructure.
+For example, in the RCP application, all positions from the same actor (output from the MOT step) can be stored in the same physical node where the PRED computation for that actor will take place when/if it arrives.
+In this example, the affinity key for all objects and tasks should be the same (the identifier of the actor, for example).
 Collocation avoids the extra overhead of fetching input data from remote processes.
-Section 4 shows evidence of this benefit: as seen in Figure 5, end-to-end latency is significantly lower and more consistent when employing correlation grouping.
-Prefetching: The knowledge encoded into correlation keys can also enable proactive decisions in anticipation of a future need.
-For example, in a multistage job a scheduler launching the first stage of a computation might also dispatch a notification that some collection of objects will be required in the second or subsequent stages.
-This is why our correlation key mechanism permits one object to be tagged with multiple keys: if we consider an ML computation such as RCP with a graphical structure, the output of one stage might be an input to multiple downstream stages, each with its own group of data dependencies.
-Using correlation keys, this would enable the platform to prefetch objects that will be needed but are not currently in cache, to transfer objects from originating node to a consuming node in anticipation of the future need, etc.
-Notice the similarity here to a publish-subscribe model, in which publishers might not know who the subscribers will be, and subscribers might not know their publishers.
-In our work, the correlation key is playing an analogous role.
-Consistency: Many edge ML applications are sensitive to data sequencing.
-Similarly to partition keys in Azure EH, correlation keys can be used to guarantee the consistency of requests.
-Requests with the same correlation key may have to be handled sequentially and in order.
-In platforms where objects sharing the same correlation key are to be stored on the same “home nodes”, one update can be used to update the entire set as a single atomic action.
-Parallelism: Requests with different correlation keys have no dependency with each other and may thus be handled in parallel.
-Furthermore, the correlation grouping mechanism enables a finer grain of parallelism control compared, for example, with a mechanism such as Azure EHs.
-In EHs, requests with different partition keys may be placed in a same partition, causing the requests to be handled sequentially, although they could be potentially have been parallelized.
-4. Scaling with Correlation Grouping
-In this section, we describe an implementation of the correlation grouping mechanism in Cascade (Song et al., 2022; Derecho Project, 2023): a state-of-the-art stream processing platform targeted at latency-sensitive ML workflows.
-Cascade was selected because it offered the highest baseline performance among open source platforms amenable to our methodology, enabling us to ask whether even lower latency and higher throughput might be feasible using an correlation grouping methodology.
-The goal of this section is to show the benefit of correlation grouping on the end-to-end (E2E) latency of the RCP application.
+Prefetching: The knowledge encoded into affinity keys can also enable proactive decisions in anticipation of a future need.
+For example, in a multistage job a scheduler launching the first stage may decide to prefetch objects that will be needed by a downstream stage, based on the affinity key of the objects and the corresponding stage, in anticipation of the computational task reaching the corresponding nodes.
+Consistency: Many edge AI applications are sensitive to data sequencing.
+Similarly to partition keys in Azure EH, affinity keys can be used to guarantee consistency.
+Objects and tasks with the same affinity key may have to be handled sequentially and in order.
+Additionally, a platform may allow objects sharing the same affinity key to be updated atomically, in a single action: this was easily achieved in our prototype, since objects with the same affinity key are stored in a same shard, as described later in Section 4.
+Parallelism: Objects and tasks with different affinity keys have no mutual dependency and may thus be handled in parallel.
+Furthermore, the affinity grouping mechanism enables a finer grain of parallelism control compared, for example, with a mechanism such as Azure EHs.
+In EHs, requests with different partition keys may be placed in the same partition, causing the requests to be handled sequentially.
+4. Evaluation on Local Cluster
+In this section, we describe an implementation of the affinity grouping mechanism in Cascade (Song et al., 2023): a state-of-the-art stream processing platform targeted at latency-sensitive AI workflows.
+Cascade was selected because it offered the highest baseline performance among open source platforms amenable to our methodology, enabling us to ask whether even lower latency and higher throughput might be feasible using an affinity grouping methodology.
+The goal of this section is to show the benefit of affinity grouping on the end-to-end (E2E) latency of the RCP application.
 We evaluate the E2E latency of the application with different workloads, while also scaling out the Cascade deployment.
@@
 4.1. RCP Application Implementation
-Our experiment used the Stanford Drone Dataset (SDD) (Robicquet et al., 2016), which contains 60 aerial videos showing campus intersections.
-The videos include several types of agents (pedestrians, cars, cyclists, skateboarders, buses, and golf carts).
-Data source “clients” stream these videos at 2.5 FPS.
+Our experiment used the Stanford Drone Dataset (SDD) (Robicquet et al., 2016), which contains aerial videos showing campus intersections.
+The videos include several types of actors (e.g.
+pedestrians, cars, cyclists).
+Data source “clients” simulate cameras by streaming these videos to the pipeline.
+The MOT and PRED steps are based on off-the-shelf AI models, and we used the source code and pre-trained models provided by the authors as much as possible: most of our modifications focused on data ingress/egress.
 All three steps are implemented in Python using the PyTorch library.
-To implement the MOT step, a multi-object tracker (Broström, 2022) that employs YOLO5 for agent detection was used.
+To implement the MOT step, a multi-object tracker (Broström, 2022) that employs YOLO5 for actor detection was used.
 StrongSORT (Du et al., 2022) and OSNet (Zhou et al., 2022b) are used for re-identification and trajectory tracking.
 We selected three videos from the SDD for training (little3, hyang5, and gates3).
-Data is uncompressed, hence each frame is approximately 8MB in size.
-State data containing positions and features of agents detected in a frame (required to re-identify agents in the following frame) ranges from a few KBs to around 10MB depending on the number of agents in the frame.
-Frames from the same video must be processed sequentially, but frames from different videos can be processed in parallel by multiple instances of the MOT step.
-Appendix A includes plots showing the number of agents detected in each frame of each video, as well as the size of the MOT state.
+Data is uncompressed, hence each frame is approximately 8MB in size (compressing and then decompressing frames is slower than just transferring them uncompressed).
+State data containing positions and features of actors detected in a frame (required to re-identify actors in the following frame) ranges from a few KBs to around 10MB depending on the number of actors in the frame.
+Frames from the same video must be processed sequentially, but frames from different videos can be processed in parallel by multiple MOT instances.
 The PRED step employs a trajectory prediction model, YNet (Mangalam et al., 2021) pretrained on the SDD data by the authors (Girase, 2022).
-PRED requires eight consecutive positions for each agent and makes no prediction if fewer than eight are available.
+PRED requires eight consecutive positions for each actor and makes no prediction if fewer than eight are available.
 Position objects are small: 10s of bytes.
-Instances of PRED for the same agent must be processed sequentially.
-Instances triggered for different agents may be executed in parallel.
-Notice that the PRED workload depends not only on how many clients are streaming frames, but also on how many agents are detected in each frame.
-CD consists of a simple algorithm that performs a linear interpolation on the predicted trajectories of identified agents, outputting a warning if any pair crosses.
+Instances of PRED for the same actor must be processed sequentially.
+Instances triggered for different actors may be executed in parallel.
+Notice that the PRED workload depends not only on how many clients are streaming frames, but also on how many actors are detected in each frame by MOT.
+CD consists of a simple algorithm that performs a linear interpolation on the predicted trajectories of identified actors, outputting a warning if any pair crosses.
 Instances of this step corresponding to the same frame and client must be processed sequentially to ensure that all pairs of trajectories will be matched with each other.
@@
 4.2. Cascade: System Architecture
-Cascade is a full-stack platform for high-speed stream processing that prioritizes low latency by hosting data and compute, avoiding copying or locking on critical data paths and leveraging acceleration technologies such as RDMA or DPDK.
-Backend components are implemented in C++, while client APIs are available in multiple languages (e.g.
-Java and Python).
-Cascade consists of a set of nodes interconnected in a complete network.
-Nodes can be clients or servers.
+Cascade (Song et al., 2023) is a full-stack platform for high-speed stream processing that prioritizes low latency by hosting data and compute, avoiding copying and locking on critical data paths and leveraging acceleration technologies such as RDMA and DPDK.
+It consists of a set of nodes (clients and servers) interconnected in a complete network, typically in an edge/cloud datacenter.
 Client nodes may issue requests to any server node.
@@
 Figure 2 shows an overview of the system architecture.
-Figure 2. Cascade architecture.
-Cascade’s storage subsystem implements a sharded Key/Value (K/V) object store.
-Server nodes are logically grouped into disjoint shards with a few nodes each.
+The storage subsystem implements a sharded Key/Value (K/V) object store.
+Server nodes are logically grouped into disjoint shards.
 Prior to our work, each object key was hashed to determine the home shard for the K/V pair.
-Cascade supports multiples levels of persistence, but our work considered only the two in-memory mode: trigger requests (which cause a task to run but leave no data behind) volatile storage requests (which replicate an object that will then be retained in memory by all members of the home shard).
-Volatile storage is limited to objects that can fit within memory.
-In the RCP scenario, for example, object sizes ranged from a few hundred bytes to a few MBs.
+Cascade supports multiples levels of persistence, but our work considered only the two in-memory mode: trigger requests, which cause a task to run but leave no data behind; and volatile storage requests, which replicate an object that will then be retained in memory by all members of the home shard.
 Cascade’s User-Defined Logic (UDL) framework is responsible for executing computational tasks.
-The code for each task is supplied by the developer as a container or in a DLL.
-In either case, a self-initialization method is expected to register one or more upcallable functions, each associated with a key prefix.
+The code for each task is supplied by the developer as a container or in a Dynamic-Link Library (DLL).
+In either case, a task is associated with a key prefix.
 On each Cascade server, an upcall will occur if that server receives a K/V pair with a matching key.
-For example, if a task is registered using the prefix “/RCP/taxis”, then an update to a K/V pair with key “/RCP/taxis/1234” would trigger that task.
-Cascade’s trigger operation will cause a single upcall to occur, on whichever node the put was sent to.
-Cascade’s volatile put has two possible behaviors, under developer control: it can upcall on every node in the shard that the key mapped to, or it can do the upcall on one randomly selected node within the shard.
-Although a single K/V trigger or volatile put will trigger a task, once the computation is running that task can access additional objects using Cascade’s K/V get API.
-Moreover, any task can issue a trigger put or a volatile put, enabling pipelined tasks in which each stage initiates the next stage.
-The potential overheads in this model are thus: (1) a P2P send operation is used to send a trigger put; (2) an atomic multicast is used to send a volatile put; (3) if a task calls get to access an object that is not homed on the node where it is executing, it must either be fetched from cache or over the network.
-Even with high-speed networking primitives supported by RDMA or DPDK, network transfers of large objects are costly, so Cascade must minimize the frequency of such events.
-Nodes have limited capacity, hence a further issue arises if the execution framework needs to enlarge the pool of compute resources.
-Lacking a warm-start mechanism, Cascade will incur the cost of fetching all needed objects.
+For example, if a task is registered using the prefix “/RCP/taxis”, then an update to a K/V pair with key “/RCP/taxis/1234” would trigger that task at whichever node the put was sent to.
+Each task can also issue another put, enabling pipelined tasks in which each stage initiates the next stage.
+In the context of this work, a potential overhead that we want to minimize occurs when a task reads an object that is not homed on the node where it is executing.
+The object then must either be fetched from cache or over the network.
+Even with high-speed networking primitives supported by RDMA or DPDK, network transfers of large objects are costly.
 Cascade supports a form of resource partitioning called an object pool.
-The basic idea is as follows: rather than treating all Cascade nodes as a single sharded service, the nodes are instead organized into groups.
+Rather than treating all Cascade nodes as a single sharded service, the nodes are instead organized into groups.
 A single server node can belong to multiple pools.
 Cascade allows the developer to configure each pool with its own properties, such as the shard size to use, and the degree of data persistence the pool will offer.
-A persistent pool can be pinned to specific storage hardware (SSD, rotating disk, Optane, etc).
-A volatile pool can be pinned to host memory, GPU memory, etc.
+For example, a volatile pool can be pinned to host memory, GPU memory, etc.
 Pools are identified by pathname prefixes: if /x/y is the prefix of a volatile pool holding GPU memory objects, /x/y/z could name a tensor residing within that GPU memory.
-4.3. Correlation Grouping on Cascade
-Most aspects of the Cascade architecture can easily accommodate the proposed correlation grouping mechanism.
-As noted earlier, data and computation requests in Cascade consist of trigger, put or get operations.
+4.3. Affinity Grouping on Cascade
+As noted earlier, requests in Cascade consist of trigger, put or get operations.
 The first two are parameterized by the key and value (uninterpreted byte vector) of the object being transmitted/saved, whereas get takes a key and returns the corresponding object.
-We implemented function $f$ as a regular expression, which is matched against request keys, but with one simplification: our initial implementation limits each request to a single correlation key  [^1] In future work we will support mappings to multiple correlation keys.
-A single key suffices for the RCP application, but multiple keys are likely to be important in more complex applications where the output of one compute stage is consumed by multiple follow-on stages, each with its own dependency set..
-The correlation key is a substring of the request key, composed of the characters that matched the regular expression.
+We implemented function $f$ as a regular expression, which is matched against request keys (the descriptor).
+The affinity key is a substring of the request key, composed of the characters that matched the regular expression.
 We employed the Hyperscan library (Wang et al., 2019) for matching regular expressions, which introduced a negligible overhead to Cascade’s critical path: according to microbenchmarks, matching the regular expressions employed in the RCP application was under 300 microseconds on average.
-To implement the application-level API of the correlation grouping mechanism, we extended the Cascade client API.
+To implement the application-level API of the affinity grouping mechanism, we extended the Cascade client API.
 The extension allows developers to register regular expressions with a specified object pool.
-Figure 3 shows a code excerpt (in Python) with an example of how object pools are created, with and without employing correlation grouping.
-The only modification made to the Cascade API was the addition of the optional argument correlation_set_regex to the method that creates an object pool (line 12).
-Any put or get operation assigns an correlation key to the corresponding object by matching the object key against the regular expression registered in the corresponding object pool (line 18).
+Listing 1 shows a code excerpt (in Python) with an example of how object pools are created, with and without employing affinity grouping.
+The only modification made to the Cascade API was the addition of the optional argument affinity_set_regex to the method that creates an object pool (line 12).
+Any put or get operation assigns an affinity key to the corresponding object by matching the object key against the regular expression registered in the corresponding object pool (line 18).
+Listing 1: Creating pools with/without affinity grouping.
 1capi = ServiceClientAPI() # Cascade client API
@@
 4
-5# creating object pool without correlation grouping
+5# creating object pool without affinity grouping
 6capi.create_object_pool("/no_grouping",
@@
 8
-9# creating object pool with correlation grouping
+9# creating object pool with affinity grouping
 10capi.create_object_pool("/grouping",
 11 subgroup_type,subgroup_index,
-12                        correlation_set_regex="_[0-9]+")
+12 affinity_set_regex="_[0-9]+")
 13
-14# putting object in the object pool that is not grouped
+14# putting in the object pool that is not grouped
 15capi.put("/no_grouping/example_1",None)
 16
-17# putting object in the object pool that is grouped
-18capi.put("/grouping/example_1",None) # correlation key is ’_1’
-Figure 3. Code excerpt showing how object pools with/without grouping are created and how objects are put/get.
-With respect to the platform-level runtime engine, we made two modifications to Cascade: (i) the mapping from key to shard within an object pool is based on the correlation key instead of the request key; and (ii) the correlation key functions and their matching expressions are registered in all nodes.
+17# putting in the object pool that is grouped
+18capi.put("/grouping/example_1",None) # affinity key is ’_1’
+With respect to the platform-level runtime engine, we made two modifications to Cascade: (i) the mapping from key to shard within an object pool is based on the affinity key instead of the request key; and (ii) the affinity key functions and their matching expressions are registered in all nodes.
 As described above, Cascade previously selected the shard that will handle a request by hashing the object key supplied with the object describing the request.
-Our modified policy selects the shard by hashing the correlation key.
-Jointly, these changes ensure that all objects with the same correlation key will be stored (and replicated) in the same shard, and route requests using that same correlation key to this shard.
-An open question concerns whether (and how) schedulers should be extended to factor in correlation grouping considerations.
-Cascade is already designed to send computations to the home nodes of needed objects, or to nodes where those objects are cached: an ideal policy from our perspective.
+Our modified policy selects the shard by hashing the affinity key.
+Jointly, these changes ensure that all objects with the same affinity key will be stored (and replicated) in the same shard, and route requests using that same affinity key to this shard.
 4.4. Computing Environment
-We now describe a series of experiments conducted on our local cluster.
-Servers in the cluster are equipped with Mellanox ConnectX-4 VPI NIC cards connected to a Mellanox SB7700 InfiniBand switch.
-This configuration results in a RDMA-capable 100Gbps network backbone.
+Servers in our local cluster are equipped with Mellanox ConnectX-4 VPI NIC cards connected to a Mellanox SB7700 InfiniBand switch, resulting in a RDMA-capable 100Gbps network backbone.
 All servers have their clocks synchronized using PTP (Eidson et al., 2002), making timestamps from different servers comparable with sub-millisecond precision.
@@
 Configuration A consists of two Intel Xeon Gold 6242 CPUs, 192 GB of memory, and an NVIDIA Tesla T4 GPU.
-This configuration is commonplace in Edge Cloud environments, where applications such as RCP are often deployed.
 Configuration B consists of two Intel Xeon E2690 v0 CPUs and 96 GB of memory, but no GPU.
-All servers run Linux kernel 5.15.0.
-Cascade nodes were deployed on Docker (version 24.0.6) containers hosted on the cluster servers.
-Containers were based on Ubuntu 22.04, with CUDA version 11.7.
+Our experiments employ up to 8 servers with configuration A, and up to 9 with B.
 Cascade was deployed with three object pools, one per step.
-Cascade nodes in pools responsible for MOT and PRED were deployed on containers hosted on servers with configuration A, since these steps require GPUs for ML inference.
-Clients and other Cascade nodes were deployed on containers hosted on servers with configuration B.
-In this section, we report results with a varying number of shards in each pool, as well as a varying number of nodes in each shard.
-Each configuration of shards/nodes is called a layout.
-Although Cascade includes a scheduler, we configured the system to place objects and trigger computations purely by correlation key hashing.
-This enables us to focus on the degree to which a purely correlation grouping placement of data and computation can improve latency.
-4.5. RCP Application Deployment on Cascade
+Cascade nodes in pools responsible for MOT and PRED were deployed on servers with configuration A, since these steps require GPUs for AI inference.
+Clients and other Cascade nodes were deployed on servers B.
+In this section, we report results with a varying number of shards in each pool.
+Each shard has always only one node (i.e.
+a physical server): increasing the size of each shard would increase replication, which is not relevant for evaluating our proposals.
+Each configuration of shards is called a layout, and we denote a layout as $x/y/z$, where values $x$, $y$, and $z$ indicate the number of shards for steps MOT, PRED, and CD, respectively.
+Although Cascade includes a scheduler, we configured the system to place objects and trigger computations purely by affinity key hashing.
+This enables us to focus on the degree to which a purely affinity grouping placement of data and computation can improve latency.
+4.5. RCP Deployment on Cascade
 Each step of the RCP application was deployed as a DLL on Cascade’s UDL framework.
-UDLs are loaded by each Cascade node on startup.
-ML models are also loaded on startup by the UDLs – all nodes have local access to models and hyperparameters.
-All required objects (MOT state data, agent positions and trajectory predictions) are stored/retrieved using Cascade K/V store.
-Each client is responsible for a different video, and sends complete video frames at a rate of 2.5 FPS.
+UDLs are loaded by each Cascade node on startup, along with models, weights, and hyperparameters.
+All required objects (MOT state data, actor positions and trajectory predictions) are stored/retrieved using Cascade K/V store.
+Each client is responsible for a different video, and sends video frames at a rate of 2.5 FPS (the rate the AI model employed in PRED was trained at).
 To send a frame, a client puts an object in the K/V store with a key that will trigger a MOT task.
-The MOT task retrieves the state data from the previous frame, and puts a separate object for each agent position found in the frame.
-Each agent position put by the MOT task triggers a PRED computation.
-Each PRED task first retrieves the past positions of the corresponding agent from the K/V store, and puts an object with the trajectory prediction.
-Each trajectory prediction triggers a CD computation, which first retrieves all trajectory predictions for the same frame available in the K/V store, and then performs the computation.
+The MOT task retrieves the state data from the previous frame, and puts a separate object for each actor position found in the frame.
+Each actor position put by the MOT task triggers a separate PRED computation.
+Each PRED task first retrieves the past positions of the corresponding actor from the K/V store, and puts an object with the trajectory prediction.
+Each trajectory prediction triggers a separate CD computation, which first retrieves all trajectory predictions for the same frame available in the K/V store, and then performs the computation.
 The result of the collision detection is put in the K/V store.
@@
 All tasks cache in memory all objects they retrieve or create.
-We employed two different workloads: single clients (little3, hyang5, and gates3), and three simultaneous clients (little3 + hyang5 + gates3).
-Each client send 700 frames, and we discard measurements from the first 100 frames.
+We employed two workloads: single clients (little3, hyang5, and gates3), and three simultaneous clients (little3 + hyang5 + gates3).
+Each client sends 700 frames, and we discard measurements from the first 100 frames.
 Each experiment is repeated three times.
-Cascade was completely cleared of objects and object pools before each run of each experiment, and a setup script was executed to create the necessary object pools.
-We grouped requests from each step using the correlation grouping mechanism in a different way.
-For MOT, grouping was based on the identifier of the client, so that all frames from a same client always went to the same shard.
-For PRED, grouping was based on agent identifiers, and for CD it was based on both the client identifier and the frame number.
-Table 1 shows all the employed object pools, examples of object keys, which computations are triggered when objects are put in the corresponding object pool (if any), the regular expressions used to group requests (if any), and the resulting correlation keys from matching the regular expressions against the example keys.
-Table 1. Object pools, example keys, triggered steps, regular expressions, and correlation keys.
-Object Pool Example Key Step Regex Correlation Key
-/collision/tracking/cameras /collision/tracking/cameras/little3_42 MOT /[a-zA-Z0-9]+_ /little3_
-/collision/tracking/states /collision/tracking/states/little3_42 /[a-zA-Z0-9]+_ /little3_
-/collision/tracking/agent_position /collision/tracking/agent_position/little3_7_42 PRED /[a-zA-Z0-9]+_[0-9]+_ /little3_7_
-/collision/prediction/agent_prediction /collision/prediction/agent_prediction/little3_42_7 CD /[a-zA-Z0-9]+_[0-9]+_ /little3_42_
-/collision/detection /collision/detection/little3_42_7_5
-Results reported in this section compare two different placement strategies, random placement and correlation grouping.
-In the random placement strategy, the RCP application is executed without using correlation grouping, i.e.
+Cascade was completely cleared of objects and object pools before each run of each experiment.
+We grouped requests from each step using the affinity grouping mechanism in a different way.
+For MOT, grouping was based on the identifier of the client, so that all frames from the same client always went to the same shard.
+For PRED, grouping was based on actor identifiers, and for CD it was based on both the client identifier and the frame number.
+Table 1 shows all the employed object pools, key examples, triggered tasks (if any), the regular expressions used to group requests (if any), and the resulting affinity keys.
+Table 1. Object pools, example keys, triggered steps, regular expressions, and affinity keys.
+Object Pool Example Key Step Regex Affinity Key
+/frames /frames/little3_42 MOT /[a-zA-Z0-9]+_ /little3_
+/states /states/little3_42 /[a-zA-Z0-9]+_ /little3_
+/positions /positions/little3_7_42 PRED /[a-zA-Z0-9]+_[0-9]+_ /little3_7_
+/predictions /predictions/little3_42_7 CD /[a-zA-Z0-9]+_[0-9]+_ /little3_42_
+/cd /cd/little3_42_7_5
+Results reported in this section compare two different placement strategies, random placement and affinity grouping.
+In the random placement strategy, the RCP application is executed without using affinity grouping, i.e.
 objects and computations are placed randomly by hashing requests keys – the standard Cascade behavior.
-In the correlation grouping strategy, we group requests as described above.
+In affinity grouping, we group requests as described above.
+Hash-based pseudo-random mapping from key to shard is standard in key-value stores.
+Our discovery is that by combining affinity grouping with randomization on the affinity key, we get the best of both worlds: excellent load-balancing and scaling with sharply higher cache hit rates, hence less data movement.
 4.6. Results
-We now report results for experiments with just one client.
-We varied the number of shards for MOT (1, 3), PRED (1, 3, 5) and CD (1, 3, 5).
-The number of nodes in each shard was 1.
+We now report results for just one client, varying the number of shards for MOT (1, 3), PRED (1, 3, 5) and CD (1, 3, 5).
 Figure 3 shows a box plot of the E2E latency (in milliseconds) for different layouts and placement strategies.
-We show only results for client gates3 for the sake of saving space.
-Results for other clients followed the same pattern.
-It is possible to see that correlation grouping significantly decreased both median and 75th percentile latencies, except for layout 1/1/1, in which the grouping mechanism had no effect since there was only 1 shard per step.
-Furthermore, increasing the number of shards did not improve latency with the random placement strategy.
-In fact, both the median and percentile latencies slightly increased as more shards were added: fetching overhead increased, due to more cache misses.
+We show only client gates3, results for other clients followed the same pattern.
+Both median and 75th percentile latencies were significantly reduced by affinity grouping, except for layout 1/1/1: grouping had no effect since there was only 1 shard per step.
+Adding shards does not help with the random placement strategy.
+Although increasing from 1/1/1 to 1/3/3 did help, increasing the number of shards even further results in higher median and percentile latencies, as fetching overheads increase due to more cache misses.
+We now increase the workload.
+Figure 4 shows the E2E latency for three simultaneous clients (little3 + hyang5 + gates3), employing different layouts and placement strategies.
+Latency was significantly lower and more consistent as the deployment scaled out when employing affinity grouping.
 Figure 3. E2E latency for gates3 on Cascade.
-We now increase the workload to three simultaneous clients.
-Figure 5 shows the E2E latency (in milliseconds) for three simultaneous clients (little3 + hyang5 + gates3), employing different layouts and placement strategies.
-Latency was significantly lower and more consistent as the deployment scaled out when employing correlation grouping.
 Figure 4. E2E latency for three clients on Cascade.
+[…]
 E2E latency for three clients on Cascade.
-Figure 5. E2E latency for three clients on Cascade.
-In order to further highlight the benefit of collocation, we disabled the object caching performed by each step of our application.
-Without caching, steps always have to fetch objects from the K/V store.
-In this context, the main difference between the random placement and correlation grouping strategies is that correlation grouping guarantees that objects being fetched are always stored locally, i.e.
-in the same Cascade node (and shard) where the computation is running.
-In the case of random placement, there is no guarantee: objects are placed randomly across shards.
-Figure 6 shows the results comparing the two placement strategies with and without the application-level caching, for a workload of three simultaneous clients, and 3/5/5 shards (MOT/PRED/CD steps, respectively), one node per shard.
-Due to the zero-copy design of Cascade, E2E latency was the same with or without caching when employing correlation grouping, since memory copies are minimized and there is no serialization overhead when an UDL makes a get request to the same node where it is running: keeping objects in the application memory or fetching them locally from the K/V store incurs virtually the same cost.
-However, in the case of random placement, disabling caching significantly increased latency and reduced throughput.
-The median E2E latency was off-scale at more than 58 seconds, hence the bar for this case is replaced with an arrow.
+To further highlight the benefit of collocation, we disabled object caching in our application.
+Each step then always has to fetch objects from the K/V store.
+Here, the main difference between random placement and affinity grouping is that the latter guarantees that objects being fetched are always stored locally, in the same Cascade node (and shard) where the task is running.
+For random placement, there is no such guarantee.
+Figure 5 shows the results for three clients and 3/5/5 shards.
+Due to the zero-copy design of Cascade, E2E latency was the same with or without caching for affinity grouping, since memory copies are minimized and there is no serialization overhead when an UDL makes a get request to the same node where it is running: keeping objects in the application memory or fetching them locally from the K/V store incurs virtually the same cost.
+However, for random placement, disabling caching significantly increased latency and reduced throughput.
+The median latency was off-scale at more than 58 seconds, hence the bar is replaced with an arrow.
 Throughput was on average 6.7 FPS, thus the pipeline was not able to handle the 7.5 FPS sent by clients collectively.
 Figure 5. E2E latency with/without caching on Cascade.
-We also evaluated the impact of replication on the E2E latency of the application.
-Different layouts with more than one node per shard were employed for this evaluation.
-On Cascade, when a shard has more than one node, any object stored in that shard is replicated to all nodes in the shard, before any computation is triggered.
-As a result, any computation triggered in such a shard will have local access to the objects in that shard, even if the previous computation (e.g.
+We also evaluated the impact of replication on E2E latency.
+Different layouts with more than one node per shard were employed.
+In Cascade, when a shard has more than one node, any object stored in that shard is replicated to all nodes in the shard, before any task is triggered.
+As a result, tasks in such a shard will have local access to the objects in that shard, even if the previous computation (e.g.
 state of the previous frame) was performed in another node.
@@
 Figure 6 shows results for different layouts.
-The first group of bars is presented as reference, and show the E2E latency for 3/5/5 shards, with one node per shard (no replication).
-The next two bars present results for 1/1/1 shards, with different number of nodes per shard (3 and 3/5/5), so there is no difference between random placement and correlation grouping.
-The last layout has 1/3/3 shards, with 2 nodes per shard, configuring a compromise between having many shards with just one node each and a single shard with many nodes.
+As a reference, the first group of bars show the latency for 3/5/5 shards, one node per shard (no replication).
+The next two bars show results for 1/1/1 shards, with different number of nodes per shard (3 and 3/5/5), so there is no difference between random placement and affinity grouping.
+The last layout has 1/3/3 shards, two nodes per shard, configuring a compromise between many shards with just one node each and a single shard with many nodes.
 Results show that replication reduced latency compared to the baseline (first bar).
-However, employing correlation grouping and multiple shards still results in better E2E latency.
-We argue that these results show the potential of further investigating the integration of correlation grouping with prefetching and scheduling.
+However, employing affinity grouping and multiple shards still results in better latency.
+We argue that these results show the potential of further investigating the integration of affinity grouping with prefetching and scheduling.
 Figure 6. E2E latency with replication on Cascade.
 4.7. Insights
-Correlation grouping significantly reduces the latency of the RCP application on Cascade, compared to other placement options Cascade is able to support.
-Additional benefits include better scalability than with random placement, where adding nodes sometimes degraded performance.
-Performance was steady across varied workloads, suggesting that correlation grouping improves hardware utilization.
-Changes to the application were primarily to register regular expressions (Table 1 during initialization, as illustrated in Figure 3).
-Support for correlation grouping required minor changes to Cascade itself, yet enabled the system to use correlation keys to place objects and route requests.
-Our results on replication indicate that there is potential to improve latency even further by integrating correlation grouping with prefetching.
-5. RCP Deployment on a Public Cloud
-A natural question to ask is whether our claim that existing SP and ML serving platforms are inadequate is correct.
+Affinity grouping significantly reduces the latency of the RCP application on Cascade, compared to other placement options.
+Although the sharding policy in Cascade makes it scalable by design, results show that affinity grouping improved Cascade scalability even further compared to the standard random placement, where adding nodes sometimes degraded performance due to increased cache misses.
+It is important to note that the scalability of affinity grouping is orthogonal to the scalability of the system as a whole.
+The scalability of the affinity grouping mechanism itself is discussed in Section 3.
+Changes to the application were primarily to register regular expressions (Table 1 during initialization, as illustrated in Listing 1).
+Support for affinity grouping required minor changes to Cascade itself, yet enabled the system to use affinity keys to place objects and route requests.
+Our results on replication indicate that there is potential to improve latency even further by integrating affinity grouping with prefetching.
+5. Evaluation on Azure Cloud
+A natural question to ask is whether our claim that existing SP and AI serving platforms are inadequate is correct.
 To address this, we now describe a second deployment of RCP on Microsoft Azure Cloud.
 Granted, Azure is a public cloud whereas Cascade primarily targets a private cluster, but we believe this is a fair comparison because it genuinely represents a state-of-the-art alternative.
-The goal of the section is to show evidence that: (i) data access patterns of applications such as RCP indeed result in extra fetching overheads; and (ii) although it is possible to overcome these overheads, it comes at the cost of a high coupling between application and deployment, since there is no unified mechanism available in modern platforms such as the proposed correlation grouping.
+The goal of the section is to show evidence that: (i) data access patterns of applications such as RCP indeed result in extra fetching overheads; and (ii) although it is possible to overcome these overheads, it comes at the cost of a high coupling between application and deployment, since there is no unified mechanism available in modern platforms such as the proposed affinity grouping.
 5.1. Deployment on Microsoft Azure Cloud
@@
 We reused the source-code of each of the 3 steps (MOT, PRED, and CD), with few modifications.
-Each of the ML models was first configured on Azure Machine Learning (AML) as real-time endpoints, using a web-services interface.
+Each of the AI models was first configured on Azure Machine Learning (AML) as real-time endpoints, using a web-services interface.
 AML is elastic: depending on load, endpoints can be backed by worker instance pools of varying size (Microsoft, 2022a).
@@
 Frames from all three videos were stored in Azure Blob Storage (Calder et al., 2011), which was also used to store MOT state data.
-Positions of agents and trajectory predictions are stored using Cosmos DB (Reagan, 2018).
+Positions of actors and trajectory predictions are stored using Cosmos DB (Reagan, 2018).
 The resulting architecture is shown in Figure 7, and is compliant with recommendations from the cloud vendor.
-The individual ML tasks were properly configured with ample resources (similar to the resources employed previously in the Cascade experiments).
+The individual AI tasks were properly configured with ample resources (similar to the resources employed previously in the Cascade experiments).
 Figure 7. RCP deployment on Azure.
-Clients stream video frames to the first SA job in the pipeline at a rate of 2.5 FPS, through an EH.
-The SA job invokes an MOT endpoint: one of the instances in the endpoint is selected by a load balancer to serve the request.
+Clients stream frames to the first SA job in the pipeline at 2.5 FPS, through an EH.
+The SA job invokes an MOT endpoint: one instance in the endpoint is selected by a load balancer to serve the request.
 The instance fetches the video frame from Blob storage and performs the inference (we assume the client has already uploaded the frame).
 The state from the previous frame is also fetched from Blob storage if it is not already cached in the instance’s memory.
-Results are sent through another EH to the next SA job responsible for invoking PRED endpoints for agent positions.
-For each invocation, the selected instance of the invoked PRED endpoint stores the new position on Cosmos DB, and also fetches the past positions of the corresponding agent that are not already cached in memory.
+Results are sent through another EH to the next SA job responsible for invoking PRED endpoints for actor positions.
+For each invocation, the selected instance of the invoked PRED endpoint stores the new position on Cosmos DB, and also fetches the past positions of the corresponding actor that are not already cached in memory.
 The CD step is performed in a similar fashion.
@@
 The partition key for MOT requests was the video name, ensuring that frames originating in any single client were processed sequentially.
-Similarly, the partition key for PRED was the agent identifier, and for CD the frame number.
+Similarly, the partition key for PRED was the actor identifier, and for CD the frame number.
 We set the batch size when invoking AML endpoints to 1, a configuration intended to minimize per-frame latency.
@@
 To vary the workload (the number of simultaneous clients streams), we first tested with a single video at a time (little3, hyang5, and gates3), then with two concurrent video streams (little3 + hyang5), and finally with three (little3 + hyang5 + gates3).
-5.2. Results: From One to Two Simultaneous Clients
+Figure 8. E2E latency for individual clients.
+Figure 9. MOT latency breakdown.
+5.2. Increasing Workload
 We now report results for experiments with just one client, a case that establishes baseline latency for each video stream, processed individually in a heavily provisioned, dedicated infrastructure.
@@
 The reason for this is that 3 instances were enough to significantly parallelize invocations to PRED.
-Increasing the number of instances even further resulted in more network overhead of fetching agent positions that offset the gain of extra parallelism, since the rate of cache misses increased.
+Increasing the number of instances even further resulted in more network overhead of fetching actor positions that offset the gain of extra parallelism, since the rate of cache misses increased.
 With 3 instances, 64ms was spent per frame fetching data in the PRED step for hyang5, while 74ms was spent on average with 5 instances.
-Figure 9. End-to-end latency for individual clients.
 We increased the workload to two simultaneous clients (little3 + hyang5), using the same configuration as the experiments above for individual clients.
@@
 The breakdown shows how much time was spent on average (in ms) on different MOT sub-steps: fetching the video frame, retrieving the state from the previous frame, inference, and uploading the state of the current frame.
-Figure 10. Latency breakdown of the MOT step.
-5.3. Results: Slashing Overheads From MOT
+Figure 10. Grouped MOT with 2 and 3 clients.
+5.3. Slashing Overheads From MOT
 Since we know that the MOT step requires state data regarding frames from the same video, we deployed a separate MOT endpoint for each video, with one instance each.
@@
 It is possible to see that latency became lower and more consistent as the number of instances for PRED and CD endpoints increased.
-Figure 11. Grouped MOT with 2 and 3 simultaneous clients.
 However, we observed that there is still potential to further cut network overheads.
-As we increased the number of PRED and CD instances, parallelism is increased, but so is the time spent fetching agent positions and trajectory predictions from Cosmos DB.
+As we increased the number of PRED and CD instances, parallelism is increased, but so is the time spent fetching actor positions and trajectory predictions from Cosmos DB.
 Figure 11 shows the latency breakdown for PRED and CD steps, with three simultaneous clients, three grouped MOT endpoints, and a varying number of PRED and CD instances (3, 5, and 7).
@@
 With 7 PRED/CD instances, for example, the average time spent fetching data per frame in PRED and CD steps was almost 200ms (roughly 100ms for each step).
-Figure 12. Latency breakdown of PRED and CD steps.
-5.4. Results: Slashing Overheads From PRED and CD
+Figure 11. PRED and CD latency breakdown.
+5.4. Slashing Overheads From PRED and CD
 Here we employ the same approach as described above for MOT, leveraging application-specific knowledge to group PRED and CD.
@@
 Thus, we manually select in the SA jobs code which endpoint to forward each request, by writing the output of the previous step to the corresponding EH.
-For PRED, this selection was based on the agent identifier, modulo the number of PRED endpoints.
+For PRED, this selection was based on the actor identifier, modulo the number of PRED endpoints.
 For CD, the selection was based on the frame number, modulo the number of CD endpoints.
-Figure 13 shows a code excerpt (simplified for the sake of presentation) from the SA job responsible for the MOT endpoint associated with the video little3.
+Listing 2 shows a code excerpt (simplified for the sake of presentation) from the SA job responsible for the MOT endpoint associated with the video little3.
 In the code, from-little3-client is an EH where the client responsible for the little3 video sends requests.
 Function mot-little3-endpoint invokes the AML endpoint containing a single instance that performs the MOT inference.
-There are two queries (lines 9 and 12) that forwards the position of each agent to a specific EH (to-pred-0 or to-pred-1) depending on the agent identifier and the number of PRED endpoints.
-The EHs to-pred-0 and to-pred-1 will deliver the agents positions to the corresponding SA jobs that will invoke their associated PRED endpoints.
+There are two queries (lines 9 and 12) that forwards the position of each actor to a specific EH (to-pred-0 or to-pred-1) depending on the actor identifier and the number of PRED endpoints.
+The EHs to-pred-0 and to-pred-1 will deliver the actors positions to the corresponding SA jobs that will invoke their associated PRED endpoints.
+Listing 2: Code excerpt for the MOT SA job.
 1WITH mot_output AS (
@@
 9SELECT * INTO [to-pred-0] FROM expanded_output
-10WHERE agent_id % num_pred_endpoints = 0
+10WHERE actor_id % num_pred_endpoints = 0
 11
 12SELECT * INTO [to-pred-1] FROM expanded_output
-13WHERE agent_id % num_pred_endpoints = 1
-Figure 13. Code excerpt from the MOT SA job for little3.
+13WHERE actor_id % num_pred_endpoints = 1
 Figure 12 shows a comparison between the pipeline with only the MOT step grouped and the pipeline with all three steps grouped.
@@
 The gain in parallelism was limited, however, we argue that more instances would be able to handle a heavier workload without an increase in latency.
-Figure 14. Grouping only MOT and grouping all steps.
+Figure 12. Grouping only MOT vs grouping all steps.
 5.5. Insights, Trade-offs, and Limitations
-Without our manual grouping, scaling in Azure is straightforward: it is only necessary to increase the number of instances in AML, streaming units in SA, and/or partitions in EHs.
+Without our manual grouping, scaling in Azure is straightforward: it suffices to increase the number of AML instances, SA streaming units and/or EH partitions.
 AML endpoints will distribute requests across all the available instances automatically according to their load.
-Furthermore, it is possible to use auto-scaling features, which will scale the deployment in/out according to predefined metrics.
-However, when employing the manual grouping strategy, scaling is done by adding or removing endpoints, which requires the application to be reconfigured.
-In the case of our deployment in Azure, changing endpoints affects client and SA jobs code, and it is also necessary to create new EHs.
-Load balancing and auto-scaling would need to be implemented by the application itself.
-Despite all trade-offs, the benefit of grouping is significant, as also observed in Section 4.
-For MOT in particular, grouping was essential to support more than one client.
-Without grouping, the overhead of fetching states caused each MOT request to take more than 400ms, which is the time it takes for each client to send the next frame request.
-This caused requests to pile up in queues at the first step of the pipeline.
-In a time-pressured scenario such as the RCP application, the 400ms threshold is of high importance: the ML model employed in the PRED step predicts agents trajectories in the next 4.8 seconds.
-If it takes too long to process each frame, due to queuing, then the prediction may not be useful any longer.
-Furthermore, since the application handles a real-time video stream, the MOT state data for each frame will only be ever used once (for the following frame).
-However, without grouping it is moved twice (upload and download), which can be considered wasteful.
-Another clear benefit of grouping is better instance utilization.
-One could argue that each video/client configures a separate application which could be deployed separately, in a dedicated infrastructure.
-However, we showed that grouping allows multiple clients to share the same infrastructure while maintaining the same level of performance, which requires less instances to be deployed since they are better utilized.
-Furthermore, such shared deployment enables new clients to be dynamically added in runtime.
-However, this would require an extra front-end step in the pipeline that decides how the grouping of a new client should be performed (we’ve done it statically for our experiments).
+Furthermore, the pattern lends itself to autoscaling.
+In contrast, with manual grouping scaling entails adding or removing endpoints, which requires that the application be reconfigured.
+In the case of Azure, changing endpoints requires code changes to client and SA jobs and the creation of new EHs, and the application itself would have to do load-balancing/auto-scaling.
+The benefit of grouping is significant, echoing our observation in Section 4.
+For MOT, grouping was essential to support more than one client.
+Without grouping, the overhead of fetching states inflates runtimes to more than 400ms, meaning that the next frame arrives before the current frame is processed (at 2.5 FPS).
+Requests will pile up at the first pipeline stage.
+For a time-pressured scenario such as the RCP application, the 400ms threshold is of high importance: the AI model employed in the PRED step predicts actors trajectories in the next 4.8 seconds.
+If it takes too long to process each frame, due to queuing, it will be too late to act.
+Furthermore, since the application handles a real-time video stream, the MOT state data for each frame will only be used once (by the subsequent frame), yet without grouping is moved twice (upload and download).
 6. Related Work
@@
 Although the authors in (Lee et al., 2018) criticize black box approaches, they focus only on caching, buffering and batching in non-pipelined applications.
-The correlation grouping mechanism can support all of these capabilities but also enables additional optimizations, such as proactive data/computation collocation and prefetching.
+The affinity grouping mechanism can support all of these capabilities but also enables additional optimizations, such as proactive data/computation collocation and prefetching.
 Existing systems offer a number of mechanisms to express application-specific knowledge and/or request grouping; these features can be found in Redis hash tag, Azure Cosmos DB, EHs and SA partition key, RDD location preferences, Apache Spark Streaming partitions, and Apache Storm stream grouping.
 However, as noted in Section 3, such mechanisms: (i) are highly coupled with deployment; (ii) do not collocate data and computation in a unified fashion; (iii) are not able to express all possible correlations between data and computation; and/or (iv) may have a non-negligible computational cost as the deployment scales out.
-The Pheromone (Yu et al., 2023) system offers a novel approach that the authors refer to as a data bucket abstraction, in which the outputs of functions are automatically grouped into buckets.
+We compare affinity grouping with these other approaches in more detail in Section 3.2.
+The Pheromone (Yu et al., 2023) system offers a data bucket abstraction, in which the outputs of functions are automatically grouped into buckets.
 Developers can then arrange for downstream functions to use data buckets as inputs.
 This enables Pheromone to place functions close to where the data is stored.
-The data bucket abstraction is similar to correlation grouping, however outputs in data buckets are always volatile (once consumed, they are garbage collected).
-The correlation grouping approach is compatible with data persistence, and we would argue that this gives greater flexibility.
-For example, some outputs may be needed between separate invocations of a pipeline (e.g.
-past positions of agents in the PRED step of the RCP application); our approach covers such cases but the data bucket model does not.
-Furthermore, the authors do not consider more complex workflows in which multiple functions may require access to the same output in different stages, or in which a function requires data located in multiple buckets.
+The data bucket abstraction is similar to affinity grouping, however outputs in data buckets are always volatile (once consumed, they are garbage collected).
+The affinity grouping approach is compatible with data persistence, and we would argue that this gives greater flexibility.
+Furthermore, the authors do not consider more complex workflows in which multiple functions may require access to the same output in different stages, or in which a function requires data from multiple buckets.
 Several works improve data locality by scheduling tasks where input data (or most of it) is located (Nishihara et al., 2017; Moritz et al., 2018; Sreekanti et al., 2020; Yu et al., 2021; Boutin et al., 2014; Crankshaw et al., 2020; Zhang et al., 2017), or by scheduling tasks that interact with each other in the same location (node, VM, or container) (Mahgoub et al., 2022; Kotni et al., 2021).
 These solutions follow a reactive approach: data is first placed without taking into account their correlations, and then tasks are placed where “most of input data” is located.
-Due to the data access patterns of latency-sensitive ML pipelines, a proactive approach, such as enabled by correlation grouping, may be necessary.
+Due to the data access patterns of latency-sensitive AI pipelines, a proactive approach, such as enabled by affinity grouping, may be necessary.
 We claim that our proposed mechanism is complementary to scheduler-based solutions.
+Finally, data collocation is widely explored in literature and thus is not a novel concept per se (Mazumdar et al., 2019).
+However, we argue that our novelty lies in providing developers with an easy-to-use mechanism that requires no knowledge of deployment/environment details, while still achieving effective collocation.
+The experiments in Appendix 5 are evidence of the limitation of modern platforms and the potential to improve latency by applying collocation principles.
+Nevertheless, achieving such collocation is challenging for developers and pose trade-offs, hence the need for a better mechanism.
+7. Beyond RCP and Cascade
+In this work we employed the RCP application as a running example, and implemented affinity grouping only on the Cascade platform.
+A natural question that arises is whether affinity grouping can be generalized to other applications and/or AI hosting platforms.
+To address such question, we discuss next how the affinity grouping mechanism could be implemented in Azure, as well as other emerging applications where affinity grouping can have a great impact.
+7.1. Affinity grouping in Azure
+In Section 4 we described an implementation on a platform based on a K/V store.
+Few modifications would be required to implement our proposals in Azure.
+For example, if Azure’s storage solutions and AML operated in a consistent fashion, data stored in Cosmos DB or Blob storage with a certain affinity key could be prefetched by physical servers running the AML instance handling requests with that affinity key.
+AML endpoints with multiple instances should ideally route requests with the same affinity key to the same instance.
+If the load balancer will add instances to reduce load on a hot-spot, it could prefetch correlated data to pre-warm those instances.
+Azure has many caching components; these could be extended to load or evict object sets, applying the identical policy objects sharing a single affinity key.
+7.2. Emerging applications
+In reinforcement learning (RL) (Sheng et al., 2025), an ML system might run for hours or days repeatedly accessing the same data objects, but not always using the same nodes for running tasks.
+A developer of such system knows which data objects are accessed by each task, and therefore can map affinity keys to tasks and data objects.
+Such mapping allows a node to fetch all needed objects for a task invocation (they all share the same affinity key) at once and in parallel, and cache them as a set (or evict them as a set).
+Such objects may be a set of cached K/V pairs, corresponding to the activation state of the training NN, needed by a Low Rank Adapter (LoRA).
+Fetching one by one incurs significantly higher delays, because it doesn’t leverage the parallelism of the network, and caching systems lack semantic knowledge of the application, which is clear for the developer and can be easily encoded through affinity grouping.
+The same rationale above also applies in the training of modern transformer-based models (e.g.
+LLMs).
+In such context, it is common practice to cache collections of Query ($Q$), Key ($K$), Value ($V$) tuples, which are a snapshot of the DNN state of some layer $L$ at step $T$, triggered by input query Q.
+Any given ($Q,K,V$) tuple is updated by only one neuron in layer $L$, but the iterated (“auto-regressive”) training algorithm requires that at each step every neuron in layer $L$ read the full set of layer $L$ tuples from the prior step, adapt its own weights (compute a gradient), apply the gradient to the layers around it (forward and back propagation), and then update its (Q, K, V) tuple.
+By hashing $Q$, $L$ and $T$ we can identify an affinity group: the entire group of tuples for this query at this time step for layer $L$ can be cached (or evicted) jointly.
+Updates will stream from individual neurons to the tuples they own, and as each step finishes, the neurons active in the next step will simply read the cached prior values.
+In contrast, suppose we viewed this affinity grouping as a key in a standard K/V store like Reddis.
+Clearly, we could form a single object collecting the set of ($Q,K,V$) tuples for layer $L$ at step $T$.
+Once serialized to a byte vector, we would have a K/V object that can be stored in a sharded key-value storage service.
+But because this one big object has contributions from all neurons in $L$, we would either need to run AllReduce simply to compute it, or pick a K/V store that supports coherent in-place updates and have each neuron read the working value, update it to add its own ($Q,K,V$) tuple, and write the result back.
+After one update per neuron, the version of the object would be complete for step $T$, and step $T+1$ could run.
+The issue here is that all the neurons in layer $L$ contribute to this object, and because they run concurrently, all contend to read, modify, and rewrite the object when creating it.
+That will take time: one layer can have many neurons, so these two alternative approaches (AllReduce or one big object holding the collection) would be very slow in comparison to an affinity-based solution, where each node only reads tuples from its own cache.
 8. Conclusion
-We proposed an correlation grouping mechanism that enables developers to express application-specific knowledge of data/computation correlations, and then for platforms to use this information to reduce latency and improve efficiency.
-Our results show that the proposed mechanism is able to maintain significantly lower and more consistent latency, as well as higher node utilization as the application workload increases and the infrastructure scales out.
-Acknowledgements.
-We are grateful to Professor Luís Rodrigues for the insightful comments on early versions of this paper.
-Partial support for our work was provided by AFRL/RY under the SWEC program, Microsoft, Nvidia and Siemens.
-Appendix A Application Workload
-Figure 15 show the workload of each video (little3, hyang5, and gates3), as well as the aggregated workload for two simultaneous videos (little3 + hyang5), and three simultaneous videos (little3 + hyang5 + gates3).
-Workload is expressed in number of agents detected in each frame.
-It is possible to see that little3 has the lowest workload across all 700 frames, while gates3 has the highest – during the first half of the videos, there are more agents in gates3 frames than in the aggregated little3 + hyang5 frames.
-Figure 15. Number of agents for different workloads.
-Figure 16 shows the size of the MOT state (in MBs) for each frame of each of the three videos employed in our experiments.
-Figure 16. Size of the MOT state for each frame of each video.
-oter>
+We proposed an affinity grouping mechanism that enables developers to express application-specific knowledge of data/computation correlations, and then for platforms to use this information to reduce latency and improve efficiency.
+Our results show that the proposed mechanism is able to maintain significantly lower latency as the application workload increases and the infrastructure scales out.
+Results also show great potential for integrating the proposed mechanism with other approaches such as prefetching and scheduling.